### v1

# Repetition Improves Language Model Embeddings

Abstract
Recent approaches to improving the extraction of text embeddings from autoregressive large language models (LLMs) have largely focused on improvements to data, backbone pretrained language models, or improving task-differentiation via instructions. In this work, we address an architectural limitation of autoregressive models: token embeddings cannot contain information from tokens that appear later in the input. To address this limitation, we propose a simple approach, “echo embeddings,” in which we repeat the input twice in context and extract embeddings from the second occurrence. We show that echo embeddings of early tokens can encode information about later tokens, allowing us to maximally leverage high-quality LLMs for embeddings. On the MTEB leaderboard, echo embeddings improve over classical embeddings by over $9\%$ zero-shot and by around $0.7\%$ when fine-tuned. Echo embeddings with a Mistral-7B model achieve state-of-the-art compared to prior open source models that do not leverage synthetic fine-tuning data. [^1] Our code and pre-trained models are released at https://github.com/jakespringer/echo-embeddings.
Repetition Improves Language Model Embeddings
Jacob Mitchell Springer  Suhas Kotha Daniel Fried  Graham Neubig  Aditi Raghunathan Carnegie Mellon University {jspringe, suhask, dfried, gneubig, aditirag}@cs.cmu.edu
1 Introduction
Neural text embeddings have a crucial role in modern approaches to information retrieval (IR), semantic similarity estimation, classification, and clustering (Ni et al., 2021b; Muennighoff et al., 2022). For example, document retrieval often leverages low-dimensional embeddings for efficient lookup: when queries and documents are encoded as vectors where semantic relationships are described by similarity in some metric space, a query lookup can be reduced to an approximate nearest-neighbor search in embedding space (Johnson et al., 2019; Vanderkam et al., 2013).
In the recent past, the dominant pretrained language model paradigm for neural embeddings have been masked language models with bidirectional attention (Ni et al., 2021a; Raffel et al., 2020; Izacard et al., 2021; Wang et al., 2022; Jiang et al., 2022; Su et al., 2022; Xiao et al., 2023a; Li et al., 2023). However, more recent literature (Ma et al., 2023; Wang et al., 2023) has begun to scale these algorithms to modern autoregressive language models such as LLaMA-2 and Mistral (Touvron et al., 2023; Jiang et al., 2023a). Developing approaches to construct embeddings from autoregressive language models is promising: for many tasks, these models are the highest quality models available (Srivastava et al., 2022).
Figure 1: Conceptual overview of echo embeddings.
In this paper, we address a striking failure mode of autoregressive language models. This failure arises from the fact that for autoregressive language models, contextualized token embeddings—the vector of last-hidden-layer activations at the position of a particular input token—do not contain information from tokens that appear later in the sentence due to the causal attention mask. We demonstrate that such embeddings can fail to appropriately determine similarity when the early tokens are superficially similar but become dissimilar in important ways when using key information from end of the input.
We propose a strategy to overcome this limitation in autoregressive models through “echo embeddings.” With this approach, we repeat the inputs so that it appears twice in the context passed to the language model, and extract embeddings from the second occurrence. Repeating the input enables the contextualized token embeddings of the second occurrence of the passage to encode information from tokens that appear later in the passage by attending to their first occurrence in the passage. We show that echo embeddings do in fact allow embeddings of the early tokens to capture information about the later tokens.
We then evaluate echo embeddings on the standard Massive Text Embedding Benchmark (MTEB) leaderboard [^2] The MTEB leaderboard can be found at https://huggingface.co/spaces/mteb/leaderboard. In the zero-shot setting, echo embeddings improve on classical embeddings by over $9\%$ and provide consistent gains across all the different tasks for a variety of language models and scale. We then perform an apples-to-apples comparison when fine-tuning embeddings from Mistral-7B and continue to see consistent gains of echo over classical across the various tasks (by $0.7\%$ on average). Strikingly, echo embeddings with the strong Mistral-7B language model allows us to achieve state-of-the-art embedding quality, enabling autoregressive language models to match prior open-source top performing models that otherwise leveraged MLMs with bidirectional attention. [^3] The contemporaneous work by Wang et al. (2023) achieves state-of-the-art accuracy on MTEB using classical embeddings from autoregressive language models, but through finetuning on high quality synthetic data, which we believe is largely orthogonal to our contribution.
The approach of echo embeddings is conceptually well-motivated, extremely simple, and generally compatible with other innovations in extracting embeddings from autoregressive language models. As language models are likely to continue to improve over the coming years, echo embeddings can be a simple but powerful twist to classical embeddings that allow us to maximally leverage autoregressive language models.
2 Preliminaries
Our goal is to extract text embeddings that map a sentence $x$ to a vector $\phi(x)\in\mathbb{R}^{d}$ such that the semantic similarity between sentences is captured as similarity between their embeddings. In practice, we use the cosine similarity between embeddings to capture semantic similarity (detailed in Appendix B).
Embeddings from language models.
We are primarily interested in the embeddings extracted from autoregressive language models, which typically have causal attention masking and are trained on a next-token objective. For brevity, we drop the term “autoregressive” in the following.
As is standard, we extract embeddings from the activations of the final hidden layer. Each input token $x_{j}$ at position $j$ is associated with a contextualized token embedding which is the hidden layer representation $\phi_{j}(x)$.
We can pool the embeddings across all the tokens in different ways. In this work, we focus on two common strategies which have been considered by prior work (Reimers and Gurevych, 2019; Muennighoff, 2022; Zhang et al., 2023a; Wang et al., 2023).
A mean token embedding over a set of indices $A$, refers to the mean contextualized token embeddings at indices in $A:\phi_{A}(x)\coloneqq\frac{1}{\left|A\right|}\sum_{t\in A}\phi_{t}(x)$.
A last-token embedding refers to the contextualized token embedding of the last token in the input sequence, written $\phi_{-1}(x)$.
Classical embeddings.
Traditionally, embeddings are computed by simply passing the sentence to the model and extracting some pooling (e.g. mean or last-token) of the contextualized embeddings corresponding to the input sentence. We will refer to embeddings created in this way as “classical embeddings”. Additionally, one might first prompt the language model with an explanation of the task of interest followed by the sentence, and then pool the contextualized embeddings of the sentence tokens like before (Su et al., 2022).
3 Echo Embeddings
In this section, we first demonstrate a failure mode of classical embeddings, and motivate a new method that we call echo embeddings that addresses this failure.
3.1 Classical Embeddings Miss Bidirectional Information.
Sentence embeddings should aggregate information across the entire sentence. However, for autoregressive language models, the contextualized embedding at position $k$ $\phi_{k}(x)$ cannot encode information about tokens $x_{k+1},x_{k+2},\ldots$. Hence, the “meaning” encoded by the embeddings of tokens at the beginning of a sentence might inaccurately suggest they are similar (or dissimilar) to other tokens without considering the influence of tokens that come later. As a simple illustration, consider the following.
$q$ $\colon{{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[She loves summer]}}}{{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [but dislikes the heat]}}}$
$s^{-}$ $\colon{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[She loves summer]}}{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [for the warm evenings]}}$
$s^{+}$ $\colon{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[Summer is her favorite]}}{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [but not the temp.]}}$
Here, the contextualized embeddings of the first half of $s^{+}$ and $s^{-}$ are both similar to $q$ because they do not attend to the second half of the sentence. As a result, the similarity between $q$ and $s^{-}$ would be overestimated by any pooling strategy that uses information from the first half. We address last-token pooling at the end of this section.
3.2 Echo Embeddings
We propose a simple fix to mitigate the failure above: we present the input sentence twice to the language model and extract contextualized embeddings from the second occurrence of the sentence. In principle, the contextualized embeddings of the second occurrence can attend to the entire sentence presented in the first occurrence. Furthermore, in order to encourage the second occurrence to actually “encode” information about the first, we instruct the language model to perform a generic task that requires using this information, e.g., “rewrite” or “repeat.”
Classical embeddings: Feed sentence $x$ to the language model and pool the contextualized embeddings of sentence $x$. Echo embeddings: Feed a prompt such as “Rewrite the sentence: $x$, rewritten sentence: xx” to the language model and pool the contextualized embeddings of the second occurence of $x$.
Key to our method is passing the sentence twice to the model and pool embeddings exclusively from the second occurrence. [^4] We find that minor variations of the echo embeddings prompt (e.g. change “rewrite” to “repeat”) work equally well and we provide an example list in Appendix B Other tricks from classical embeddings such as prompting the model with the downstream task of interest can be applied to echo embeddings as well.
Figure 2: (left) We take the echo embeddings of only AA for query $q$ and sentences $s^{-},s^{+}$ and plot the distribution of cosine similarities, showing that echo embeddings encode later information in earlier tokens. (right) We plot the accuracy of classical and echo embeddings when the sentences have similar beginnings (Structure 1) and dissimilar beginnings (Structure 2).
3.3 Repetition Captures Bidirectional Info
In the previous section, we argued that classical embeddings suffer from the issue that contextualized embeddings of early tokens can miss out on information from the later tokens. But can simply repetition via echo embeddings solve this issue? We aim to test this by extracting embeddings from Mistral-7B on a simple controlled synthetic setting.
Given a query $q\colon[A,B]$ we construct sentence pairs $s^{+},s^{-}$ as follows. We make the first parts of each sentence identical to the query, but differ only in their second parts,
$q\colon[A,B]$; $s^{+}\colon[A,B^{+}]$; $s^{-}\colon[A,B^{-}]$,
where $B^{+}$ is $B$ but paraphrased and $B^{-}$ is semantically dissimilar to $B$. We query GPT-4 to generate examples of this structure. We describe the full procedure and prompts in Appendix B.
With classical embeddings, the contextualized embeddings of $A$ parts of $s^{+},s^{-},q$ are identical by construction. To test whether echo embeddings can meaningfully distinguish $s^{+}$ and $s^{-}$ despite having identical initial tokens, we take the mean over just the $A$-portion of the echo embeddings and plot the cosine similarities $\operatorname{Sim}(q,s^{+})$ and $\operatorname{Sim}(q,s^{-})$ in Figure 2 (left). We find that $\operatorname{Sim}(q,s^{+})$ is typically larger than $\operatorname{Sim}(q,s^{-})$. Since we are only pooling the echo embeddings of the A-portion, any distinction between $s^{+},s^{-}$ must come from the echo embeddings of $A$ capturing information from the later parts of the sentence. This showcases that current autoregressive language models can in fact allow early tokens to capture information from later tokens via echo embeddings.
3.4 Classical vs. Echo on Synthetic Data
In Section 3.3 we demonstrated that echo embeddings encode bidirectional information. However, is this sufficient to recover from the failure mode of classical embeddings? Further, where will we expect echo embeddings to improve over classical embeddings? Here, we compare echo and classical embeddings on synthetic data to answer both of these questions.
Datasets.
We sample datasets according to two structures depending on whether the discriminating information between $s^{+}$ and $s^{-}$ is in the first half (structure S1) or second half (structure S2) of the sentence. Using the structures below, we generate samples using GPT-4, as in the previous section (full details in the appendix):
(S1) $q\colon[A,B]$, $s^{+}\colon[A^{+},B^{+}]$, $s^{-}\colon[A^{+},B^{-}]$
(S2) $q\colon[A,B]$, $s^{+}\colon[A^{+},B^{+}]$, $s^{-}\colon[A^{-},B^{+}]$.
We measure the accuracy of identifying which of two sentences $s^{+}$ and $s^{-}$ is closer to the query as measured by the cosine similarity in the embeddings. We compare classical vs echo embeddings when using mean pooling to aggregate embeddings extracted from Mistral-7B model. We use mean token embedding rather than last token because last token embeddings can be quite fragile in a zero-shot setting (Section 5.1).
Results.
We present results on the two different structures in Figure 2 (right). We see that classical embeddings struggle on Structure 1—when the distinguishing information is at the beginning, the embeddings corresponding to these early tokens exaggerate similarity between $q$ and $s^{-}$ because they do not encode the information provided by $B^{-}$. In contrast, echo embeddings are able to successfully determine the more similar sentence, presumably because embeddings of the $A$-portion now also encode information about the later parts (demonstrated in Section 3.3). As a control for other reasons echo embeddings outperform classical embeddings, we also compare these embeddings on structure two, where early tokens provide discriminative signal without needing the later context. As expected, both classical and echo embeddings achieve good performance in this setting. All in all, this analysis on synthetic data demonstrates that zero-shot classical embeddings do not encode information about later context in early token embeddings, but echo embeddings can do so.
Does last-token pooling resolve the failure of classical embeddings?
The embedding of the last token $\phi_{-1}(x)$ is, in principle, can encode information from the entire input. However, we posit that the last-token pooling strategy is highly brittle and can depend too strongly on the tokens near the end of the input. To verify this, we compare the accuracy of mean token pooling and last-token pooling for classical and echo embeddings in two settings. First, we evaluate on the original synthetic data of Structure 1 (Figure 3, left). Second, we evaluate on the same data, but where we append a uniformly randomly selected token to the end of each example (Figure 3, right). While last-token pooling has high accuracy on the original toy data (though still lower accuracy than echo embeddings), it fails to perform well on the noisy examples. Echo embeddings with mean token pooling, however, are robust to the noise.
While this particular distribution of noise is artificial, it highlight that last-token pooling can be sensitive to noise in the last token. We verify in Section 5.1 that last-token embeddings perform poorly on real data. Thus, even if last-token embeddings address the inability of mean token classical embeddings to encode information from tokens that appear later in the sequence, they are not practical due to their sensitivity to noise.
Does last-token pooling resolve the failure after finetuning?
In practice, it is common to finetune embeddings on a sentence similarity objective. It is hard to delineate the degree to which this failure mode remains after finetuning. Nonetheless, we demonstrate in Section 5.2 that our method improves in the finetuning setting, even when using last-token pooling.
Figure 3: We compare the accuracy of classical and echo embeddings, and mean and last-token pooling on sentences which have similar beginnings (Structure 1). We plot these accuracies on the original data (left), and the data in which a single uniformly randomly chosen token is appended to the end of each sentence (right).
4 Methodology
In Section 3, we explored how echo embeddings can improve over classical embeddings by addressing a fundamental failure mode. In this section, we describe the methodology by which we evaluate echo embeddings on large scale real datasets in both the zero-shot and finetuning settings. While finetuning is currently necessary to achieve state-of-the-art performance, zero-shot embeddings have the advantage that they do not require expensive finetuning on top of a pretrained language model. Zero-shot results can also more clearly show how different embedding strategies work on real datasets.
4.1 Constructing Zero-shot Embeddings.
We extract zero-shot embeddings via different strategies from three language models: Mistral-7B, LLaMA-2-7B, and LLaMA-2-13B. We select the instruction-finetuned model for each of them. Refer to Appendix A.2 for additional information on the base models. Recent literature suggests that the performance of language models on zero-shot tasks can be highly variable depending on the exact wording and template of the prompts (Sclar et al., 2023). Thus, for each of the embedding strategies we consider, we perform prompt randomization where we sample prompts by randomizing the exact wording, punctuation, and capitalization of the prompt. We describe the sampling process and the exact prompts that we use in Appendix C.
Baselines.
We compare our proposed echo embeddings (Section 3.2) to classical embeddings and two additional baselines:
  • Last-token embeddings: We mentioned in the Section 3 that last-token embeddings tend to underperform in comparsion to mean token embeddings, and thus we compare on real data.
  • Summarization: We also compare zero-shot embeddings obtained via the strategy proposed by Jiang et al. (2023b). Here, they instruct the model to summarize the input in a single word and then take the last token embedding $\phi_{-1}(x)$ as the pooled embedding of the sentence.
4.2 Constructing Finetuned Embeddings.
We adopt the conventional sentence embedding training setup (Reimers and Gurevych, 2019) where we train with a contrastive learning objective to encourage the embeddings of similar text to be close. We extract embeddings in a slightly different fashion compared to the zero-shot setting above in order to keep the finetuning methodology as similar as possible to the existing literature.
Extracting embeddings.
The training and evaluation data is separated into two categories: symmetric data, in which sentences are drawn from a single distribution (such as for sentence similarity), and asymmetric, in which the data consists of both queries and documents (such as for retrieval). We adopt a separate prompt for symmetric inputs and queries, and for documents. We construct classical embeddings by encoding text $S$ using the following prompts:
Queries & Symm. Documents
Instruct: {instruction}
$\text{Query: }S$
$\text{Document: }S$
For echo embeddings, we use the prompts, where $S$ represents the input and $S^{\prime}=S$:
Queries & Symm. Documents
Instruct: {instruction}
$\text{Query: }S$ $\text{Query again: }S^{\prime}$
$\text{Document: }S$ $\text{Document again: }S^{\prime}$
In this case, {instruction} refers to the task instruction, which specifies a description of the task that the embedding will be used for. We adopt the instructions from Wang et al. (2023), and provide a list of the instructions in Appendix D. We append an end-of-sentence token to the end of each input, and we allow the input embedding of this token to be trainable.
Datasets.
We train on a collection of publicly available datasets that encompass both symmetric and asymmetric data that are standard training datasets in the embedding literature. We list and describe each of the datasets in Appendix D.
Optimization.
To finetune the model, we optimize the SimCSE loss with in-batch and mined hard negatives. Since this is standard, we defer discussion of this to Appendix D. Each batch is constructed by sampling a dataset from our set of training dataset, and then collecting examples from only this dataset. We use GradCache to train with a large batch size (2048) with limited GPU memory (Gao et al., 2021a). We train with LoRA instead of full finetuning, with $r=16$ and $\alpha=16$. We choose $\tau=1/50$ and a learning rate of $8\times 10^{-4}$. We use the Mistral-7B instruction-tuned model as a backbone (Jiang et al., 2023a). Our choices aim to be consistent with prior literature (Wang et al., 2023; Su et al., 2022; Zhang et al., 2023a).
4.3 Massive Text Embedding Benchmark
For evaluation, we use the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022). For this paper, we focus on the English-language subset of the benchmark. MTEB is composed of a collection of 56 datasets that are grouped into different embedding tasks: classification, clustering, pair classification, reranking, retrieval, sentence similarity (STS), and summarization. The goal is to construct general purpose embeddings that are useful for solving each of the tasks. More information about MTEB is specified in Appendix A.1.
For the fine-tuning setting, we evaluate on the entire English-language subset. In the zero-shot setting, for convenience, we only evaluate on a subset of MTEB. We describe this subset in Appendix D.
5 Experiments
Strategy Model Pool  Clas. P. Cls. Clus. Retr. STS Rera. Average
Main results:
Echo (ours) Mistral 7B Mean   64.06 75.26 27.02 23.61 72.40 60.00 55.07
Classical Mistral 7B Mean   58.21 73.87 23.85 20.35 56.97 54.44 45.88
Prior work:
Summarization Mistral 7B Last 66.01 81.82 26.48 19.13 70.13 66.24 54.96
Ablations:
Echo Mistral 7B Last   63.11 57.93 12.82 2.97 39.14 47.35 36.60
Classical Mistral 7B Last   58.23 46.64 13.51 2.60 33.97 46.51 32.52
Echo LLaMA 7B Mean   61.64 66.29 25.11 16.12 66.18 56.35 50.26
Classical LLaMA 7B Mean   56.61 68.46 23.22 18.63 56.49 53.26 44.65
Echo LLaMA 13B Mean   64.65 74.57 25.72 26.58 72.20 62.68 55.60
Classical LLaMA 13B Mean   58.50 65.06 24.22 18.92 57.47 56.38 45.15
Table 1: Zero-shot scores on MTEB tasks for Mistral-7B. We use a retrieval validation set (FiQA2018) to select the best prompt. Refer to Appendix C for the scores with alternative validation sets. Top: Comparison of echo embeddings to classical embeddings. Center: Summarization approach to constructing embeddings (Jiang et al., 2023b). Bottom: Ablations, including last-token pooling and LLaMA-2-{7B, 13B}.
5.1 Evaluation of Zero-shot Embeddings
We compare the performance of classical, echo, and summarization embeddings on MTEB tasks (Table 1). We validate using a retrieval dataset from MTEB (FiQA2018) as a validation set as described in Section 4.3. We report the scores using alternative validation sets in Appendix C.
Echo embeddings outperform classical embeddings zero-shot.
We see that echo embeddings outperform classical embeddings by a large margin: on average, by nearly 10 points for Mistral-7B. Further, this performance increase is consistent across every MTEB category, across models (LLaMA-2 vs Mistral), and across scale (7B vs 13B). This demonstrates that echo embeddings can significantly improve the performance of embeddings on real data, suggesting that the failure mode of classical embeddings that we describe in Section 3 can affect performance on real data.
Qualitative comparison of classical and echo embeddings.
In Section 3, we demonstrate that classical embeddings overestimate the similarity between examples which are superficially similar based on tokens that appear early in the sequence. To build intuition that this applies to realistic data, we present the sentence pair from STSBenchmark, a sentence similarity task from MTEB, in which echo embeddings reduce error the most:
[…]
which has a ground-truth score of 0 (out of $5$) similarity. The sentence pair for which echo embeddings reduces error the least is:
[…]
which has a ground-truth similarity of $2$ (out of $5$). We provide more examples in the Appendix Table 7.
For this example, notice, that the sentence pair $(x_{1},x_{2})$ on which echo embeddings improve error the most has exactly the property we identify as a failure mode for classical embeddings: the sentence is superficially similar for the first few tokens. On the other hand $(y_{1},y_{2})$ does not have this property.
Quantitative evaluation of the failure mode.
The above example builds intuition that, even on real data, classical embeddings fail to properly estimate similarity on examples which are superficially similar in the early tokens. We quantitatively measure the degree to which classical and echo embeddings fail on sentences which are similar for early tokens, and for sentences which are not. We find that classical embeddings systematically fail on examples which exhibit this structure, while echo embeddings do not. For convenience, we defer the discussion of these experiments and the results to Appendix C.1.
Last-token vs mean token pooling.
We find that last-token embeddings are substantially worse than mean token embeddings in the zero-shot setting, despite the fact that in principle, the last token in the sequence can encode information from all other tokens. In practice, it is clear that the last token does not encode sufficient information to achieve strong performance on MTEB in the zero shot setting.
Echo embeddings vs summarization.
We find that the average performance across the tested MTEB datasets is similar between echo and classical embeddings. Summarization does encourage the last token in the sequence to encode information about the entire sentence. We find that summarization is much more sensitive to the exact prompt while echo embeddings are robust to such minor variations (see Figure 5 in Appendix C). We suspect that echo embeddings are more robust as a result of more directly trying to encode bidirectional information into the embeddings.
5.2 Evaluation of Finetuned Embeddings
Strategy Model Pool   Clas. Clus. P. Cls. Rera. Retr. STS Average
Main results:
Echo (ours) Mistral 7B Last 77.43 46.32 87.34 58.14 55.52 82.56 64.68
Classical Mistral 7B Last  76.57 45.78 86.37 56.71 54.87 82.03 63.98
Prior work:
UAE-Large-V1 (MLM)  75.58 46.73 87.25 59.88 54.66 84.54 64.64
multilingual-e5-large (MLM)  77.56 47.10 86.19 58.58 52.47 84.78 64.41
bge-large-en-v1.5 (MLM)  75.97 46.08 87.12 60.03 54.29 83.11 64.23
udever-bloom-7b (autoregr.)  72.13 40.81 85.4 55.91 49.34 83.01 60.63
sgpt-5.8b (autoregr.)  68.13 40.34 82.00 56.56 50.25 78.10 58.93
e5-mistral-7b555e5-mistral-7b was recently released and leverages high quality synthetic data to achieve strong performance which is not publicly released. We report their performance, but we do not explicitly compare to them (Wang et al., 2023). (autoregr.)  78.47 50.26 88.34 60.21 56.89 84.63 66.63
Ablations:
Echo Mistral 7B Mean  77.00 44.94 87.73 58.30 55.11 82.52 64.22
Classical Mistral 7B Mean  76.26 42.68 86.31 57.58 53.75 81.53 62.96
Classical Mistral 7B-bidir. Last  76.70 45.94 88.15 57.23 54.96 82.42 64.23
Table 2: Finetuning scores on MTEB tasks. Top: Apples-to-apples comparison of echo embeddings and classical embeddings in which we use echo embeddings and classical embeddings with last-token pooling, with the same training setup. Center: Performance of recent open source embedding models, annotated by base model type, masked-language model or autoregressive. Bottom: Ablations for finetuning: using mean token embeddings (first two lines) and using a bidirectional architecture (last line).
Different embeddings on the MTEB leaderboard are often fine-tuned on different datasets. In order to perform an apples-to-apples comparison between embedding strategies, we fine-tune both echo and classical embeddings on the exact same datasets (described in Section 4.2). We report the results in Table 2. This table also includes a comparison to prior state-of-the-art methods using masked language models (MLM) and autoregressive language models. Further, we evaluate a number of ablations to determine the role of pooling strategy and architecture.
Echo embeddings outperform classical embeddings after finetuning.
We observe that echo embeddings consistently outperform classical embeddings on each category even after finetuning. Hence, the fundamental gap we find between classical and echo embeddings in Section 3 and in our zero-shot experiments persist after fine-tuning.
Comparison to prior state-of-the-art models.
We present comparisons to both prior MLM-based embeddings and prior autoregressive-language-model embeddings, listing the open-source models from the MTEB leaderboard. It is striking that MLMs vastly outperformed autoregressive models until recently. Our classical embeddings outperform the previous-best autoregressive language model. This is a result of using the strongest public 7B parameter language model (Mistral) and more fine-tuning data. However, despite these choices, classical embeddings do not outperform prior MLM-based approaches, perhaps because MLMs encode bidirectional context unlike classical embeddings from autoregressive models. Interestingly, echo embeddings allow us to close the gap to achieve state-of-the-art (on average) with an autoregressive model compared to prior open-sourced models on the leaderboard that used MLMs. A recent exception is the concurrent work by Wang et al. (2023) that use synthetic data to improve classical embeddings extracted from Mistral-7B. Their synthetic data is not publicly available, but the apples-to-apples comparison between classical and echo embeddings we performed suggests that echo embeddings could provide gains further gains over the numbers reported in (Wang et al., 2023) when fine-tuning with synthetic data.
Why doesn’t last-token pooling close the gap?
Since classical last-token embeddings can attend to every other token, they do not necessarily suffer from the failure mode that we highlighted in Section 3. Last token does not reliably capture relevant information in a zero-shot setting, but this could have been bridged via fine-tuning. It is thus surprising that, even after finetuning last-token embeddings that could (in principle) encode any embedding function, echo embeddings outperform classical embeddings. We identify two hypotheses that may explain this performance gap: (1) While last-token embeddings can attend to every token, the intermediate representations of earlier tokens cannot. If last-token pooling derives information from the internal representations of earlier tokens, by attending to these representations, last-token classical embeddings may still suffer from the failure mode of the earlier tokens. (2) If the post-finetuning performance benefits from the model initialization point, last-token classical embeddings may suffer: in Section 5.1 we show that last-token embeddings achieve poor zero-shot performance. We leave it to future work to explore these hypotheses. We do, however, observe that the gap between last- and mean token echo embeddings is smaller than the gap between last- and mean token classical embeddings, suggesting that echo embeddings can especially improve the quality of mean token embeddings.
Can we relax autoregressive language models to a bidirectional architecture and fine-tune?
To test the role of architecture, we finetune Mistral-7B on the same setup described in Section 4.2 but modified the architecture so as to remove the causal attention mask. While the initial weights are identical to Mistral-7B, this new model has bidirectional attention. We observe that the performance of bidirectional classical embeddings are better than our standard (causal) classical embeddings, but worse than echo embeddings. This suggests that the architecture alone is not sufficient to improve performance.
6 Related Work
Sentence embeddings.
Dense low-dimensional vectors representing textual semantics has been widely studied and applied. Early approaches involved computing embeddings for individual words (Hinton, 1984; Rumelhart et al., 1986; Elman, 1990; Mikolov et al., 2013; Pennington et al., 2014). Later work aims to compute dense vectors representing the semantics of entire sequences by combining or composing word vectors (Le and Mikolov, 2014; Iyyer et al., 2015; Kiros et al., 2015; Socher et al., 2011; Tai et al., 2015; Wang et al., 2016; Wieting et al., 2015). Khattab and Zaharia (2020) propose to use late interaction between document and query vectors to improve retrieval performance. Reimers and Gurevych (2019) propose S-BERT which takes a pretrained BERT (Devlin et al., 2018) and trains with a triplet loss on anchor sentences, semantically similar positive examples, and semantically dissimilar negative examples. More recent approaches typically adopt this approach with different pretrained models and a contrastive objective such as InfoNCE (Oord et al., 2018) or SimCSE (Gao et al., 2021b). Ni et al. (2021a) with Ni et al. (2021b) extend this approach to the T5 architecture (Raffel et al., 2020). Multiple papers use an additional unsupervised contrastive objective (Izacard et al., 2021; Wang et al., 2022). Other papers propose including prompts to improve task-specific embedding performance (Jiang et al., 2022; Su et al., 2022). Some work combines multiple of these training objectives and approaches (Xiao et al., 2023a; Li et al., 2023). Notably, except for the most recent approaches, nearly all embeddings were based upon bidirectional architectures that were often pretrained with a masked-language modeling objective.
Next-token language modeling for embeddings.
A series of papers aim to construct high quality embeddings from autoregressive large language models. Multiple papers apply the fine-tuning approach of S-BERT to language models but using a trained GPT (Radford et al., 2018) as the backbone architecture (Muennighoff, 2022; Zhang et al., 2023a). Ma et al. (2023) adopts this approach but for LLaMA-2 (Touvron et al., 2023). Jiang et al. (2023b) extracts embeddings by asking a language model to summarize the input sentence. Wang et al. (2023) is concurrent to our work and improves embeddings by adding synthetic training data and trains on Mistral (Jiang et al., 2023a).
Zero-shot embeddings.
Most recent sentence embeddings research has focused on improving finetuning. Reimers and Gurevych (2019) demonstrates that without finetuning, BERT has low-quality embeddings. To our knowledge, Jiang et al. (2023b) is the only paper that constructs zero-shot embeddings for autoregressive language models.
7 Conclusion
We have compared classical and echo embeddings in a toy example, on real data in the zero-shot setting, and after finetuning. With the toy data, we identified a failure mode of autoregressive classical embeddings, which we have shown can be recovered with echo embeddings. Our result motivates the development of higher quality embeddings which are important in retrieval applications.
In addition, until recently, masked language models largely dominated the MTEB leaderboard, despite often having an order of magnitude fewer parameters, having been trained on substantially less data, and performing worse on other benchmarks of interest to the natural language processing community. While our results do not explicitly explain the surprising success of masked language models, they do suggest that next-token language models suffer from an inherent drawback that may have stifled their performance until they became performant enough to compensate for this shortcoming. We believe that our embedding strategy achieves the best of both worlds: we gain the capability of next-token language models while recovering from the failure mode that next-token language models do not encode information about future tokens in their contextualized token embeddings.
8 Limitations
Despite the success of echo embeddings, the method has limitations. First, while echo embeddings achieve superior performance to classical embeddings, they require double the inference cost to pass two copies of the input sequence to the model. Though this is double the training cost for a fixed number of training steps, we show in Appendix D that echo embeddings achieve improved performance even when matching compute. Second, we do not fully explain why echo embeddings are improved in comparison to classical embeddings after finetuning even though there is no representational limitation. We leave it to future work to understand the exact underlying mechanisms for this improvement.
Acknowledgements
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE2140739. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
This research was supported by the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
This work was supported in part by the AI2050 program at Schmidt Sciences (Grant #G2264481).
We gratefully acknowledge the support of Apple.
Appendix A General Information for Reproducibility
In this section we include information that might aid in reproducibility that is not specific to any specific setting in the paper.
A.1 Massive Text Embedding Benchmark
The Massive Text Embedding Benchmark (MTEB) is a collection of datasets from seven categories: classification, clustering, pair classification, reranking, retrieval, sentence similarity (STS), and summarization. The leaderboard is published at https://huggingface.co/spaces/mteb/leaderboard. The list of datasets and their descriptions can be found at Muennighoff et al. (2022) in Appendix A.
A.2 Base Model HuggingFace IDs
In this paper, we use the following models:
Appendix B Echo Embeddings: Additional Information
In this section, we aim to describe the additional details that were omitted from Section 2 and 3.
Cosine Similarity.
As discussed in Section 2, we often use the cosine similarity to measure the similarity in embeddings. Recall that given two sentences $x$ and $y$, we wish to determine the degree to which they are semantically similar. Cosine similarity,
$\operatorname{Sim}(x,y)\coloneqq\frac{\left\langle\phi(x),\phi(y)\right\rangle}{\|\phi(x)\|\|\phi(y)\|},$ (1)
measures the similarity between the embeddings of $x$ and $y$ for any embedding function $\phi\colon\mathcal{X}\to R^{d}$. The cosine similarity is used for our experiments in Sections 3, and as the similarity function for training in 5. All MTEB datasets use cosine similarity to compute similarity with the exception of the classification datasets, in which similarity is not explicitly measured, and the clustering datasets, which use Euclidean distance,
$\operatorname{Sim}(x,y)\coloneqq\|\phi(x)-\phi(y)\|,$ (2)
as a metric.
Prompts for Section 3.
For these experiments, we only evaluate with a single prompting strategy. For classical embeddings, we encode a sentence $S$ using the prompt:
$x=\text{Write a sentence: $S$}$
We take the pooled embedding to be the mean token embedding $\phi_{S}(x)$. For echo embeddings, we encode a sentence $S$ using the prompt:
$x=\begin{array}[]{ll}\text{Rewrite the following sentence: $S$}\\ \text{The rewritten sentence: $S^{\prime}$}\end{array}$
where $S^{\prime}=S$ and we let our pooled embedding be the mean token embedding $\phi_{S^{\prime}}(x)$. We do not evaluate with the last-token pooling strategy in this Section.
General Prompting Guidelines.
Throughout the paper, we use a variety of different prompts to construct embeddings. In Section C, we demonstrate that for zero-shot embeddings, the exact wording or template used as a prompting strategy does not have a strong effect on the performance of MTEB tasks, with the exception of for the summarization approach. This implies, in general, that classical embeddings and echo embeddings should be robust to the exact choice of prompts. The important component of echo embeddings is instead the structure: the input text should be repeated twice when computing embeddings, and the embeddings should be taken over the second occurrence of the input text.
Example classical embedding structures:
$\text{Say the sentence: }S$
$\text{Write the phrase: }S$
$\text{Complete the query: }S$
$\text{Explain the text: }S$
Example echo embedding structures:
$\text{Repeat the sentence: }S$ $\text{The sentence again: }S^{\prime}$
$\text{Rephrase the query: }S$ $\text{The query rephrased: }S^{\prime}$
$\text{Fill in the blank: }S$ $\text{The blanks filled in: }S^{\prime}$
$\text{Rewrite the text: }S$ $\text{The sentence rewritten: }S^{\prime}$
Toy data.
We provide a subset of the toy data from Section 2. For Structure 1, the data is given in Table 4. For Structure 2, the data is given in Table 5. For Structure 3, the data is given in Table 6. In all cases, the data is generated by GPT4. The data from Structure 1 is generated from the following GPT4 prompt, and the other structures are generated from minor variations on this:
Together, we need to generate sentence triplets. Each triplet will have the following form:
- sentence 1 can be anything, be creative here.
- sentence 2 must represent something opposite to sentence 1, however, it is important that the first half of the sentence is exactly the same as the first half of sentence 2. The only difference in wording can be in the second half of the sentence.
- sentence 3 should be extremely similar to sentence 1 and semantically equivalent, but slightly re-worded
Here is an example:
"sentence1": "I like to eat apples and bananas but I really hate almost every other fruit.",
"sentence2": "I like to eat apples and bananas and I also enjoy also every other fruit",
"sentence3": "I like two fruits: apples and bananas but I hate nearly all fruits other than these.",
The first half of the sentence should be relatively short, less than 10 words, but the second half should be long, at least 10 words. Give more examples, and write them in json format. Be creative!
Appendix C Additional Zero-shot Results
In this section, we describe the omitted methodology and results for the zero-shot section.
Prompt sampling procedure.
Here we describe the prompt sampling procedure and then provide the prompts that we use for the zero shot:
  1. 1.
    Choose an instruction. For classical embeddings, we choose from {Write, Say, Complete, Explain}. For echo embeddings, we choose from {Repeat, Rewrite, Rephrase, Fill in the blank}. For summarization, we choose from {Summarize, Categorize, Understand, Analyze}.
  2. 2.
    Choose a wording for the instruction. For example, if we chose “Say” as the instruction, then we would choose from {Say a sentence, Say a paragraph, Say something, Say a response, Say a query, Say a prompt}. For summarization, we also choose a second part of the wording, as the summarization strategy requires that the summary be in one word: {in one word, with a single word, succinctly with one word, in a unique one-word way, in a single word, in a word}.
  3. 3.
    Choose a separator, which include colons, commas, newlines.
  4. 4.
    Choose a prefix, which includes markers to indicate the first and appearance of the input.
  5. 5.
    Classical prompts have the form: “{instruction} {separator} {prefix} SS”.
  6. 6.
[…]
  7. 7.
    Summarization prompts have the form: “{instruction0} {separator} {prefix} $S$ {instruction1} {separator}”.
For classical, we choose the prompts:
Write a sentence I] S
Write a prompt!
(I) S
Write some text
PROMPT-S
SAY A PARAGRAPH | SENTENCE 0] S
Say a query QUERY: S
Say a sentence!
[A] S
COMPLETE THE PROMPT Text (1) S
Complete the query SENTENCE 0) S
Complete the sentence:-S
Explain a query text 0 S
Explain a prompt | Sentence 1> S
EXPLAIN A SENTENCE Prompt (1) S
For echo, we choose the prompts:
Repeat The Paragraph.
query 1) S.
query 2) S
Repeat the response.
1) S.
AGAIN 2) S
REPEAT THE SENTENCE :: PROMPT
S :: RESPONSE
S
Rewrite the query | QUERY (A) S | (B) S
Rewrite the text. SENTENCE A) S. B) S
Rewrite the response | query A] S | query B] S
Rephrase the sentence:@S:Again@S
Rephrase The Sentence!
Text <> S!
Answer <> S
REPHRASE THE QUERY Sentence a) S Answer b) S
Fill in the blank in the prompt:
Query a) S:
Query b) S
FILL IN THE BLANK IN THE RESPONSE | Sentence A) S | Sentence B) S
Fill in the blank in the paragraph.
Text | S.
Response | S
For summarization, we use the prompts:
SUMMARIZE THE QUERY.
Prompt: SIN A WORD.
Summarize the sentence!
PROMPT <1> SSuccinctly With One Word!
SUMMARIZE THE PARAGRAPH. PROMPT (0) SIN A WORD.
CATEGORIZE THE PROMPT query
SWith a single word
Categorize the query | prompt [1] Sin a word |
CATEGORIZE THE SENTENCE.
Prompt <1> SIN A WORD.
Understand the sentence
@SIn a single word
Understand The Prompt:QUERY [0] Sin a single word:}
UNDERSTAND THE PARAGRAPH:Text I] SSuccinctly with one word:}
Analyze the sentence.
Sentence SIn A Unique One-word Way.
Analyze the response! query a> SIN A UNIQUE ONE-WORD WAY!
Analyze The Prompt
Sentence a> SIn a unique one-word way
Subset of MTEB for zero-shot evaluation.
We evalaute on the following subset of MTEB: FiQA2018, SCIDOCS, SciFact, NFCorpus, TwitterSemEval2015, TwitterURLCorpus, ImdbClassification, AmazonReviewsClassification, TweetSentimentExtractionClassification, MTOPDomainClassification, TwentyNewsgroupsClustering, BiorxivClusteringS2S, MedrxivClusteringS2S, StackOverflowDupQuestions, AskUbuntuDupQuestions, SciDocsRR, BIOSSES, STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark, and SICK-R.
Measuring the sensitivity of different embedding strategies to prompting.
We plot the sensitivity of classical, repetition, and summarization to different choices of prompts for different models in Figures 56, and 7. We also extend to plotting on all tested datasets individually in Figures 89, and 10. We observe that summarization is highly sensitive to the exact prompt used. However, neither classical nor echo were particularly sensitive in any case. Consistently, mean token pooling outperformed last token pooling by a large factor.
Evaluation of zero-shot results with different validation sets.
We include the zero results of validation using different MTEB datasets. For validation, we select one dataset from each category, as follows: classification: ImdbClassification; Pair Classification: TwitterSemEval2015; Clustering: TwentyNewsgroupsClustering; Retrieval: FiQA2018; STS: STSBenchmark, Reranking: StackOverflowDupQuestions. We plot these results for different models in Tables 89, and 10. We observe similar results across different validation sets, with minor variations in the performance. In addition, we the performance of each dataset when the prompts have been validated with FiQA2018 in Tables 1112, and 13.
Figure 4: We plot the histogram distribution of the difference between the predicted rank and the ground truth rank of sentence pairs in STS datasets. When predicted rank is larger than the ground truth rank—when the rank difference is positive—then the embedding has overestimated the similarity of this pair. Similarly, negative values imply that the the rank is underestimated. We plot the distribution of these ranks for both classical and echo embeddings where we split the data into two groups: one in which sentences are similar in the first part of the sentence (top 10% by first-half similarity), and another in which sentences are similar in the second part of the sentence (top 10% by second-half similarity).
C.1 Validating the connection between our synthetic data experiments and real data.
In Section 3, we hypothesized that classical embeddings would overestimate similarity on sentences where the first half of the sentence are similar, and that echo embeddings would recover from this failure mode. In order to test this hypothesis, we exact a set of examples from the STS datasets included in the MTEB benchmark in which the first half of the sentence is similar, and measure the degree to which the similarity is overestimated.
As a control, we also select points which are similar in the second half of the sentence, and measure the degree to which similarity is overestimated. By comparing the degree to which sentences which are similar in the first half are overestimated in similarity, and the degree to which sentences which are similar in the second half are overestimated, then we can identify if classical embeddings overestimate similarity in specifically sentences which are similar in the first half. Thus, under our hypothesis, we expect that, for classical embeddings, sentences which are similar in the first half are overestimated in similarity more than sentences that are similar in their second half. On the other hand, we expect that, for echo embeddings, the degree to which similarity is over- or underestimated is independent of whether the sentences are similar in the first or second half of the sentence.
Identifying examples based on similarity in the first/second part of the sentence.
We aim to determine which sentences are most similar in the first half of the sentence or in the second half of the sentence. For each sentence pair $x,y$, we split the sentences in half by number of words, yielding $x=[x_{1},x_{2}]$, and $y=[y_{1},y_{2}]$. We select sentences which are most similar in the first half by using the off-the-shelf masked-language-model-based embedding model bge-base-en-v1.5 (Xiao et al., 2023b). To select sentences that are similar in the first half, we measure the cosine similarity $\operatorname{Sim}(x_{1},y_{1})$ and take the top 10% of sentence pairs $x,y$ which have the highest cosine similarity. Similarly, to select sentences which are similar in the second half, we collect the top 10% of examples by $\operatorname{Sim}(x_{2},y_{2})$. We collect examples from each of the STS datasets in MTEB.
Measuring sentence similarity estimation error.
We must determine the degree to which classical and echo embeddings overestimate similarity. The STS datasets contain sentences pairs which are ranked by similarity: the sentences which are most similar have the highest ground-truth ranking, and the least similar sentences have the lowest. We will denote the ranking of sentence pair $i$ as $r_{i}$. We compute an estimated ranking $\{\hat{r}_{i}\}$ by ranking sentence pairs by the cosine similarity between their embeddings. We can compare the error in our estimated ranking by taking the rank difference $\operatorname{Err}_{i}=\hat{r}_{i}-r_{i}$. When $\operatorname{Err}_{i}>0$, we say that the $i$th sentence pair is overestimated in similarity, and similarly underestimated when $\operatorname{Err}_{i}<0$.
Results.
We plot the the distribution over rank differences for sentences which are similar in the first half and sentences which are similar in the second half for echo and classical embeddings, from all STS datasets (Figure 4). We also highlight the means of the distributions. In accordance with our hypothesis, we observe that for classical embeddings, sentences which are similar in the first half are generally overestimated in similarity more than sentences which are similar in the second half of the sentence, suggesting that classical embeddings fail particularly on sentences that are similar in early tokens. Further, we generally observe no difference between the estimation error distributions for echo embeddings, which demonstrates that echo embeddings recover from this particular failure mode.
There are some notable counterexamples: BIOSSES does not exhibit this trend, but has few examples and thus the results may arise from noise alone. Further, STS22 exhibits identical distributions in estimation error between sentences which are similar in the first half and sentences which are similar in the second half, for both classical and echo embeddings. It is unclear why this trend fails to hold for STS22. Nonetheless, the trend holds for every other dataset, suggesting that the conceptual failure of classical embeddings that we identified in Section 3 generalizes to real data.
Qualitative examples.
In addition, we provide qualitative examples of sentence pairs from STSBenchmark where echo embeddings reduce error most, and where echo embeddings reduce error least, in comparison to classical embeddings. More precisely, we plot the top and bottom 7 examples ranked by $|\operatorname{Err}^{\text{classical}}_{i}|-|\operatorname{Err}^{\text{echo}}_{i}|$, where $\operatorname{Err}^{\text{classical}}_{i}$ represents the rank difference of the $i$th example of classical embeddings, and $\operatorname{Err}^{\text{echo}}_{i}$ is similar but for echo embeddings (Table 7).
Appendix D Additional Finetuning Results
In this section, we address the omitted details from the finetuning results of the main paper.
Training Datasets.
We follow the setup of Wang et al. (2023), and use the following datasets: ELI5 (sample ratio 0.1) (Fan et al., 2019), HotpotQA (Yang et al., 2018), FEVER (Thorne et al., 2018), MIRACL (Zhang et al., 2023b), MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) (Bajaj et al., 2018), NQ (Karpukhin et al., 2020), NLI (Gao et al., 2021b), SQuAD (Karpukhin et al., 2020), TriviaQA (Karpukhin et al., 2020), Quora Duplicate Questions (sample ratio 0.1) (DataCanary et al., 2017), Mr- TyDi (Zhang et al., 2021), DuReader (Qiu et al., 2022), and T2Ranking (sample ratio 0.5) (Xie et al., 2023). We use approximately 1.5M training examples.
GPUs.
Training a model takes approximately two days on 4 A100 GPUs.
Instructions for finetuning datasets.
We also follow the setup of Wang et al. (2023), and use the instructions in Table 3. For evaluation, we use the instructions found in Table 14.
Models on MTEB leaderboard.
We compare our implementation of classical and echo embeddings to state-of-the-art approaches on MTEB. Namely, we display results for UAE-Large-V1 (Li and Li, 2023), multilingual-e5-large (Wang et al., 2024), bge-large-en-v1.5 (Xiao et al., 2023b), udever-bloom-7b (Zhang et al., 2023a), sgpt-5.8b (Muennighoff, 2022), e5-mistral-7b (concurrent work) (Wang et al., 2023).
Additional ablations.
We plot additional ablations, including ablating the role of instructions during training and evaluation, as well as providing an evaluation at step 280 (out of 720 total steps), which is approximately $1/3$ of the duration of training (Table 15). We note that echo embeddings still outperform classical embeddings in this setting.
Performance over training time.
We plot the performance over the duration of training for a subset of MTEB tasks in Figure 11. Surprisingly, task performance decreases over training for many tasks.
Computational benefits of echo embeddings.
From Table 15, we observe that even after approximately $1/3$ of the total training duration (less than $1/2$), echo embeddings achieve performance higher than classical embeddings achieve after an entire epoch (Table 2). Echo embeddings requires twice the computational cost of classical embeddings. However, this result suggests that despite this additional cost per embedding, training with echo embeddings can save on training costs by requiring less than half an epoch of training to outperform classical embeddings. Further, since each data point is only seen once, it implies that echo embeddings are much more data efficient than classical embeddings, which may be helpful when data is costly or difficult to acquire.
All results.
We plot the results for every MTEB dataset for echo embeddings, for classical embeddings, and for bidirectional embeddings in Table 16.
NLI Given a premise, retrieve a hypothesis that is entailed by the premise
NLI Retrieve semantically similar text
DuReader Given a Chinese search query, retrieve web passages that answer the question
ELI5 Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
FEVER Given a claim, retrieve documents that support or refute the claim
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
MIRACL Given a question, retrieve Wikipedia passages that answer the question
MrTyDi Given a question, retrieve Wikipedia passages that answer the question
MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query
MSMARCO Document Given a web search query, retrieve relevant documents that answer the query
NQ Given a question, retrieve Wikipedia passages that answer the question
QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question
QuoraDuplicates Find questions that have the same meaning as the input question
Squad Retrieve Wikipedia passages that answer the question
T2Ranking Given a Chinese search query, retrieve web passages that answer the question
TriviaQA Retrieve Wikipedia passages that answer the question
Table 3: Instructions for finetuning datasets.
Training objective.
For the training objective, we use the SimCSE loss (Gao et al., 2021b). It is defined,
$\ell_{i}=-\log\frac{\exp\left(\operatorname{Sim}\left(h_{i},h_{i}^{+}\right)/\tau\right)}{\sum_{j=1}^{N}\exp(\operatorname{Sim}\left(h_{i},h_{j}^{-}\right)/\tau)}.$ (3)
In this loss function, $h_{i}$ represents a query (or a reference sentence when the data is symmetric), $h_{i}^{+}$ represents a positive example associated with $h_{i}$, and $\{h_{j}^{-}\}_{j=1}^{N}$ represents the set of negatives associated with the example, including mined hard negatives.
$q$ $^{-} s+s^{+}$
She loves to travel in summer, especially to cold destinations, avoiding hot and crowded places She loves to travel in summer, but prefers to visit hot and bustling tourist spots In summer, she adores traveling, specifically to chilly locations, steering clear of warm, populous areas
The cat often sits by the window, dreaming of chasing birds and enjoying the warm sunshine The cat often sits by the window, but is too lazy to dream of chasing anything Frequently, the cat lounges near the window, imagining bird pursuits and basking in the sunlight
He reads books every night, finding solace in fiction and escaping from the stresses of daily life He reads books every night, yet he feels that non-fiction is more engaging and informative Nightly, he immerses himself in books, seeking comfort in stories and evading everyday tensions
They play music loudly in the evening, filling their home with energetic beats and vibrant melodies They play music loudly in the evening, but only soothing classical tunes to relax In the evenings, they blast tunes, their house resonating with lively rhythms and bright harmonies
She paints landscapes on weekends, expressing her creativity through vibrant colors and abstract forms She paints landscapes on weekends, preferring realistic and detailed depictions of nature On weekends, she engages in landscape painting, showcasing her artistic flair with lively hues and unconventional shapes
The children eagerly await winter, dreaming of snowball fights and building snowmen The children eagerly await winter, yet they dislike the cold and prefer staying indoors During winter, the kids are excited, imagining snow battles and constructing snow figures
He often jokes at parties, becoming the center of attention with his witty humor He often jokes at parties, but tends to alienate others with his sarcasm At social gatherings, he frequently makes jokes, captivating the crowd with his clever wit
She collects antique vases, adoring their unique designs and historical significance She collects antique vases, but is indifferent to their history and focuses on their resale value Her hobby is gathering old vases, cherishing their distinct patterns and the stories they hold
The band plays rock music loudly, thrilling audiences with energetic performances and powerful lyrics The band plays rock music loudly, but often receives complaints for being too noisy Performing rock loudly, the band excites crowds with dynamic shows and impactful words
He prefers working at night, enjoying the quiet and focusing better without distractions He prefers working at night, despite feeling more tired and less productive Nighttime is his preferred work period, appreciating the tranquility and concentrated environment
She writes poetry in her free time, pouring her emotions and experiences into each verse She writes poetry in her free time, but struggles to find inspiration and motivation During her leisure, she crafts poems, infusing her feelings and life stories into every line
Table 4: Examples of Structure 1 from Section 3
$q$ $^{-} s+s^{+}$
On sunny days, I often find myself longing for the cool breeze of the ocean and the sound of waves crashing, as I enjoy outdoor activities During rainy days, I usually prefer the warmth and quiet of my home, as I enjoy outdoor activities When the sun is shining, I tend to crave the refreshing sea air and the rhythmic sound of the ocean, since I relish spending time outdoors
As a lover of classical music, I spend hours listening to Beethoven and Bach, reveling in the complexity of their compositions, though I’m fond of playing the guitar Despite my preference for rock music, I rarely spend time on music other than playing my favorite tunes on the guitar, though I’m fond of playing the guitar Being an enthusiast of classical melodies, I often indulge in lengthy sessions of Beethoven and Bach, appreciating the intricacies of their work, as I delight in guitar playing
In the world of literature, I have an insatiable appetite for mystery novels and spend countless evenings unraveling their plots, but I adore reading poetry Contrary to my usual tastes, I rarely delve into mystery novels and prefer lighter reading materials, but I adore reading poetry As a fervent reader, my passion lies in the twists and turns of mystery stories, which I often explore during long nights, yet I cherish reading poetry
Growing up in a bustling city, I’ve always been surrounded by the constant hum of activity and the bright city lights, which makes me appreciate quiet countryside walks Having been raised in a tranquil rural area, I’m more accustomed to the sounds of nature and open fields, which makes me appreciate quiet countryside walks Raised in the lively atmosphere of an urban environment, I’m used to the never-ending city noise and glowing nights, leading me to enjoy the serenity of rural strolls
Ever since I was a child, fascinated by the vastness of the universe, I would spend countless nights gazing at the stars through my telescope, dreaming of exploring distant galaxies, yet I still find solace in simple nature hikes Though I’ve always been more interested in the immediate world around me, preferring to focus on the tangible and the present, I rarely look up at the night sky, yet I still find solace in simple nature hikes From my early years, captivated by the infinity of space, I devoted many nights to star-gazing and imagining interstellar journeys, but I also enjoy the peace of nature walks
Growing up with a passion for culinary arts, experimenting with exotic ingredients and complex recipes, and often spending whole days in the kitchen perfecting new dishes, I also have a deep appreciation for classic literature Despite my lack of interest in cooking and a preference for simple, quick meals that require minimal preparation, I’m not one to spend time in the kitchen, I also have a deep appreciation for classic literature Since childhood, I’ve been enthusiastic about cooking, often trying out unusual ingredients and intricate recipes, dedicating entire days to refining my culinary creations, and I equally cherish classic literary works
Table 5: Examples of Structure 2 from Section 3
$q$ $^{-} s+s^{+}$
SShe loves to travel in summer, especially to cold destinations, avoiding hot and crowded places She loves to travel in summer, but prefers to visit hot and bustling tourist spots She loves to travel in summer, specifically to chilly locations, steering clear of warm, populous areas
The cat often sits by the window, dreaming of chasing birds and enjoying the warm sunshine The cat often sits by the window, but is too lazy to dream of chasing anything The cat often sits by the window, imagining bird pursuits and basking in the sunlight
He reads books every night, finding solace in fiction and escaping from the stresses of daily life He reads books every night, yet he feels that non-fiction is more engaging and informative He reads books every night, seeking comfort in stories and evading everyday tensions
They play music loudly in the evening, filling their home with energetic beats and vibrant melodies They play music loudly in the evening, but only soothing classical tunes to relax They play music loudly in the evening, their house resonating with lively rhythms and bright harmonies
She paints landscapes on weekends, expressing her creativity through vibrant colors and abstract forms She paints landscapes on weekends, preferring realistic and detailed depictions of nature She paints landscapes on weekends, showcasing her artistic flair with lively hues and unconventional shapes
The children eagerly await winter, dreaming of snowball fights and building snowmen The children eagerly await winter, yet they dislike the cold and prefer staying indoors The children eagerly await winter, imagining snow battles and constructing snow figures
He often jokes at parties, becoming the center of attention with his witty humor He often jokes at parties, but tends to alienate others with his sarcasm He often jokes at parties, captivating the crowd with his clever wit
She collects antique vases, adoring their unique designs and historical significance She collects antique vases, but is indifferent to their history and focuses on their resale value She collects antique vases, cherishing their distinct patterns and the stories they hold
The band plays rock music loudly, thrilling audiences with energetic performances and powerful lyrics The band plays rock music loudly, but often receives complaints for being too noisy The band plays rock music loudly, the band excites crowds with dynamic shows and impactful words
He prefers working at night, enjoying the quiet and focusing better without distractions He prefers working at night, despite feeling more tired and less productive He prefers working at night, appreciating the tranquility and concentrated environment
She writes poetry in her free time, pouring her emotions and experiences into each verse She writes poetry in her free time, but struggles to find inspiration and motivation She writes poetry in her free time, infusing her feelings and life stories into every line
Table 6: Examples of Structure 3 from Section 3
Most improved Least improved
Sentence 1 Sentence 2 Score Sentence 1 Sentence 2 Score
The best thing you can do is to know your stuff. The best thing to do is to overcome the fussiness. 0.0 Sometime if you really want it you might need to pay an agency to get the place for you. You could probably get a tour agency to do it for you but it would cost you. 2.0
It really doesn’t matter. It doesn’t matter unless it is really far off. 3.0 There are three options: There are only three options: 5.0
I think it’s fine to ask this question. I think it is okay to ask the question. 5.0 Bremer said one initiative is to launch a US\$70 million nationwide program in the next two weeks to clean up neighborhoods and build community projects. Bremer said he would launch a \$70-million program in the next two weeks to clean up neighborhoods across Iraq and build community projects, but gave no details. 3.6
What kind of insulation is it? What kind of floors are above? 0.0 "Tony’s not feeling well," Spurs coach Gregg Popovich said. We’re thrilled to be up 3-2,” Coach Gregg Popovich said Wednesday. 1.6
It depends entirely on your company and your contract. I guess it depends on the nature of your contract. 4.0 Shares of Mandalay closed down eight cents to \$29.42, before the earnings were announced. Shares of Mandalay closed down 8 cents at \$29.42 Thursday. 4.0
You need to read a lot to know what you like and what you don’t. You have to know what you want to do. 0.0 Singapore reported no suspected SARS cases Wednesday, but officials quarantined 70 people who had contact with the Taiwanese patient. Still, Singapore quarantined 70 people who had been in close contact with the scientist. 3.0
I would say you can do it, but it wouldn’t be advised. Personally, I would say not unless it suits you. 2.0 The dollar was at 117.85 yen against the Japanese currency, up 0.1 percent. Against the Swiss franc the dollar was at 1.3289 francs, up 0.5 percent on the day. 1.333
Table 7: Example sentences from STSBenchmark in which zero-shot echo embeddings with Mistral 7B most improve (left) and least improve (right).
Figure 5: Variance over different prompting strategies for zero-shot Mistral-7B.
Figure 6: Variance over different prompting strategies for zero-shot LLaMa-2-7B.
Figure 7: Variance over different prompting strategies for zero-shot LLaMa-2-13B.
Figure 8: Variance over different prompting strategies for all evaluated datasets for zero-shot Mistral-7B.
Figure 9: Variance over different prompting strategies for all evaluated datasets for zero-shot LLaMa-2-7B.
Figure 10: Variance over different prompting strategies for all evaluated datasets for zero-shot LLaMa-2-13B.
Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
Classical
Classification 59.20 73.80 24.16 20.57 58.59 54.54 46.79
Pair Classification 58.73 71.40 24.32 20.39 59.00 54.42 46.64
Clustering 58.23 72.62 23.90 18.64 56.68 54.82 45.37
Retrieval 58.21 73.87 23.85 20.35 56.97 54.44 45.88
STS 58.31 44.03 13.07 2.63 38.95 46.77 34.63
Reranking 58.13 71.77 24.20 20.23 58.59 54.89 46.43
Echo
Classification 64.50 74.65 25.93 22.52 73.81 59.41 55.57
Pair Classification 64.15 75.93 22.25 18.35 72.75 58.47 54.15
Clustering 61.54 71.04 26.32 15.88 68.18 60.27 51.81
Retrieval 64.06 75.26 27.02 23.61 72.40 60.00 55.07
STS 64.50 74.65 25.93 22.52 73.81 59.41 55.57
Reranking 64.15 75.93 22.25 18.35 72.75 58.47 54.15
Summarization
Classification 66.62 78.95 21.79 14.68 72.13 64.24 55.22
Pair Classification 66.62 78.95 21.79 14.68 72.13 64.24 55.22
Clustering 66.66 79.59 28.08 11.88 67.30 65.19 53.43
Retrieval 66.01 81.82 26.48 19.13 70.13 66.24 54.96
STS 66.01 81.82 26.48 19.13 70.13 66.24 54.96
Reranking 63.19 75.22 26.09 20.52 65.98 59.05 51.55
Table 8: Scores for additional zero-shot validation datasets on Mistral-7B.
Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
Classical
Classification 57.59 68.65 23.72 18.06 57.19 54.59 45.14
Pair Classification 57.56 70.18 23.51 18.54 58.24 54.40 45.79
Clustering 57.14 69.91 23.35 16.98 57.66 55.38 45.25
Retrieval 56.61 68.46 23.22 18.63 56.49 53.26 44.65
STS 57.56 70.18 23.51 18.54 58.24 54.40 45.79
Reranking 56.65 66.54 22.46 10.48 55.97 54.44 42.98
Echo
Classification 62.24 67.96 23.60 14.33 65.79 55.44 49.85
Pair Classification 63.42 72.52 21.11 17.35 68.16 54.98 51.47
Clustering 60.12 66.74 23.45 11.60 64.45 56.31 48.75
Retrieval 61.64 66.29 25.11 16.12 66.18 56.35 50.26
STS 63.15 68.74 23.65 16.38 69.37 57.75 51.96
Reranking 62.30 74.23 24.69 18.17 65.07 56.76 50.51
Summarization
Classification 63.96 77.93 21.89 15.93 67.07 63.39 52.34
Pair Classification 63.96 77.93 21.89 15.93 67.07 63.39 52.34
Clustering 61.60 69.47 24.44 5.28 57.53 57.62 45.85
Retrieval 64.90 78.74 26.63 15.59 70.15 65.43 54.02
STS 64.90 78.74 26.63 15.59 70.15 65.43 54.02
Reranking 60.54 69.73 26.40 15.82 61.60 58.80 47.83
Table 9: Scores for additional zero-shot validation datasets on LLaMa-2-7B.
Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
Classical
Classification 58.24 71.65 23.91 21.79 58.74 56.37 46.66
Pair Classification 58.10 73.30 23.01 16.97 57.83 56.17 45.52
Clustering 58.61 67.47 23.30 15.51 57.93 56.86 45.05
Retrieval 58.50 65.06 24.22 18.92 57.47 56.38 45.15
STS 58.24 71.65 23.91 21.79 58.74 56.37 46.66
Reranking 58.61 67.47 23.30 15.51 57.93 56.86 45.05
Echo
Classification 64.15 74.22 25.02 27.58 70.81 61.43 55.02
Pair Classification 64.57 77.63 22.56 24.08 69.85 59.89 53.55
Clustering 63.26 73.50 25.10 27.48 69.04 61.81 54.32
Retrieval 64.65 74.57 25.72 26.58 72.20 62.68 55.60
STS 63.16 75.98 24.08 27.56 71.00 61.84 54.85
Reranking 62.90 70.58 25.53 22.11 68.82 62.38 53.02
Summarization
Classification 66.02 79.06 26.47 22.20 67.91 64.90 54.52
Pair Classification 66.02 79.06 26.47 22.20 67.91 64.90 54.52
Clustering 63.84 71.98 21.99 7.48 56.96 59.41 46.50
Retrieval 66.02 79.06 26.47 22.20 67.91 64.90 54.52
STS 66.02 79.06 26.47 22.20 67.91 64.90 54.52
Reranking 61.19 69.63 26.38 19.62 60.76 62.79 48.36
Table 10: Scores for additional zero-shot validation datasets on LLaMa-2-13B.
Dataset Classical Echo Summarization
FiQA2018 (retrieval) 7.89 12.74 12.43
SCIDOCS (retrieval) 3.60 4.88 9.97
SciFact (retrieval) 45.39 49.36 29.90
NFCorpus (retrieval) 12.07 16.57 17.51
TwitterSemEval20. (pair_classification) 47.81 62.49 59.79
TwitterURLCorpus (pair_classification) 73.87 75.26 81.82
ImdbClassificati. (classification) 72.50 72.02 82.78
AmazonReviewsCla. (classification) 37.09 40.72 45.58
TweetSentimentEx. (classification) 53.70 58.76 61.74
MTOPDomainClassi. (classification) 83.85 92.71 90.72
TwentyNewsgroups. (clustering) 20.84 29.48 30.11
BiorxivClusterin. (clustering) 23.47 27.61 27.21
MedrxivClusterin. (clustering) 24.23 26.42 25.75
StackOverflowDup. (reranking) 35.85 42.71 40.32
AskUbuntuDupQues. (reranking) 49.49 54.09 57.17
SciDocsRR (reranking) 59.38 65.91 75.30
BIOSSES (sts) 59.05 78.19 66.06
STS12 (sts) 42.01 58.43 64.62
STS13 (sts) 59.66 78.53 78.45
STS14 (sts) 50.69 68.42 71.00
STS15 (sts) 61.81 78.82 78.29
STS16 (sts) 57.03 77.52 77.40
STS17 (sts) 68.08 82.14 78.80
STS22 (sts) 61.23 57.60 47.07
STSBenchmark (sts) 47.55 73.85 77.39
SICK-R (sts) 53.19 71.95 69.48
Average 45.88 55.07 54.96
Table 11: Evaluation of all MTEB datasets for zero-shot for Mistral-7B.
Dataset Classical Echo Summarization
FiQA2018 (retrieval) 6.48 12.38 9.00
SCIDOCS (retrieval) 3.72 4.38 8.33
SciFact (retrieval) 42.18 30.61 23.01
NFCorpus (retrieval) 10.01 13.38 15.43
TwitterSemEval20. (pair_classification) 44.11 54.66 54.27
TwitterURLCorpus (pair_classification) 68.46 66.29 78.74
ImdbClassificati. (classification) 71.65 73.11 85.83
AmazonReviewsCla. (classification) 36.16 40.68 44.77
TweetSentimentEx. (classification) 52.04 54.85 59.96
MTOPDomainClassi. (classification) 81.63 89.38 89.97
TwentyNewsgroups. (clustering) 15.88 23.42 32.28
BiorxivClusterin. (clustering) 23.13 25.92 27.79
MedrxivClusterin. (clustering) 23.31 24.30 25.48
StackOverflowDup. (reranking) 35.57 40.82 35.63
AskUbuntuDupQues. (reranking) 48.51 51.42 56.09
SciDocsRR (reranking) 58.01 61.29 74.76
BIOSSES (sts) 65.31 71.96 68.04
STS12 (sts) 41.84 52.40 60.20
STS13 (sts) 58.43 72.40 76.31
STS14 (sts) 49.21 61.24 68.73
STS15 (sts) 60.03 72.67 75.59
STS16 (sts) 56.40 73.51 76.71
STS17 (sts) 62.31 71.87 79.38
STS22 (sts) 59.48 55.21 55.69
STSBenchmark (sts) 49.45 65.73 76.42
SICK-R (sts) 55.35 64.39 70.69
Average 44.65 50.26 54.02
Table 12: Evaluation of all MTEB datasets for zero-shot for LLaMa-2-7B.
Dataset Classical Echo Summarization
FiQA2018 (retrieval) 8.31 18.07 9.43
SCIDOCS (retrieval) 4.87 7.56 10.38
SciFact (retrieval) 41.64 50.55 40.19
NFCorpus (retrieval) 10.26 21.63 16.02
TwitterSemEval20. (pair_classification) 42.43 62.85 59.55
TwitterURLCorpus (pair_classification) 65.06 74.57 79.06
ImdbClassificati. (classification) 71.82 75.44 91.86
AmazonReviewsCla. (classification) 37.88 43.25 50.60
TweetSentimentEx. (classification) 52.95 58.18 59.93
MTOPDomainClassi. (classification) 84.67 92.52 87.51
TwentyNewsgroups. (clustering) 17.21 25.98 32.08
BiorxivClusterin. (clustering) 24.95 26.75 28.30
MedrxivClusterin. (clustering) 23.49 24.70 24.64
StackOverflowDup. (reranking) 37.24 44.86 38.44
AskUbuntuDupQues. (reranking) 50.74 55.21 54.15
SciDocsRR (reranking) 62.03 70.15 75.65
BIOSSES (sts) 63.26 77.60 69.33
STS12 (sts) 51.80 59.36 51.17
STS13 (sts) 61.59 79.01 76.08
STS14 (sts) 49.69 69.75 66.62
STS15 (sts) 58.48 79.86 73.75
STS16 (sts) 53.18 76.75 77.40
STS17 (sts) 65.10 80.41 75.88
STS22 (sts) 59.00 56.84 49.23
STSBenchmark (sts) 44.80 71.31 75.17
SICK-R (sts) 55.13 70.27 71.70
Average 45.15 55.60 54.52
Table 13: Evaluation of all MTEB datasets for zero-shot for LLaMa-2-13B.
AmazonCounterfactualCls. Classify a given Amazon customer review text as either counterfactual or not counterfactual
AmazonPolarityCls. Classify Amazon reviews into positive or negative sentiment
AmazonReviewsCls. Classify the given Amazon review into its appropriate rating category
Banking77Cls. Given a online banking query, find the corresponding intents
EmotionCls. Classify the emotion expressed in the given Twitter message into one of the six emotions: anger, fear, joy, love, sadness, and surprise
ImdbCls. Classify the sentiment expressed in the given movie review text from the IMDB dataset
MassiveIntentCls. Given a user utterance as query, find the user intents
MassiveScenarioCls. Given a user utterance as query, find the user scenarios
MTOPDomainCls. Classify the intent domain of the given utterance in task-oriented conversation
MTOPIntentCls. Classify the intent of the given utterance in task-oriented conversation
ToxicConversationsCls. Classify the given comments as either toxic or not toxic
TweetSentimentExtractionCls. Classify the sentiment of a given tweet as either positive, negative, or neutral
ArxivClusteringP2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts
ArxivClusteringS2S Identify the main and secondary category of Arxiv papers based on the titles
BiorxivClusteringP2P Identify the main category of Biorxiv papers based on the titles and abstracts
BiorxivClusteringS2S Identify the main category of Biorxiv papers based on the titles
MedrxivClusteringP2P Identify the main category of Medrxiv papers based on the titles and abstracts
MedrxivClusteringS2S Identify the main category of Medrxiv papers based on the titles
RedditClustering Identify the topic or theme of Reddit posts based on the titles
RedditClusteringP2P Identify the topic or theme of Reddit posts based on the titles and posts
StackExchangeClustering Identify the topic or theme of StackExchange posts based on the titles
StackExchangeClusteringP2P Identify the topic or theme of StackExchange posts based on the given paragraphs
TwentyNewsgroupsClustering Identify the topic or theme of the given news articles
SprintDuplicateQuestions Retrieve duplicate questions from Sprint forum
TwitterSemEval2015 Retrieve tweets that are semantically similar to the given tweet
TwitterURLCorpus Retrieve tweets that are semantically similar to the given tweet
AskUbuntuDupQuestions Retrieve duplicate questions from AskUbuntu forum
MindSmallReranking Retrieve relevant news articles based on user browsing history
SciDocsRR Given a title of a scientific paper, retrieve the titles of other relevant papers
StackOverflowDupQuestions Retrieve duplicate questions from StackOverflow forum
ArguAna Given a claim, find documents that refute the claim
ClimateFEVER Given a claim about climate change, retrieve documents that support or refute the claim
CQADupstackAndroidRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackEnglishRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackGamingRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackGisRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackMathematicaRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackPhysicsRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackProgrammersRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackStatsRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackTexRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackUnixRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackWebmastersRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
CQADupstackWordpressRetr. Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
DBPedia Given a query, retrieve relevant entity descriptions from DBPedia
FEVER Given a claim, retrieve documents that support or refute the claim
FiQA2018 Given a financial question, retrieve user replies that best answer the question
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
MSMARCO Given a web search query, retrieve relevant passages that answer the query
NFCorpus Given a question, retrieve relevant documents that best answer the question
NQ Given a question, retrieve Wikipedia passages that answer the question
QuoraRetr. Given a question, retrieve questions that are semantically equivalent to the given question
SCIDOCS Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
SciFact Given a scientific claim, retrieve documents that support or refute the claim
Touche2020 Given a question, retrieve detailed and persuasive arguments that answer the question
TRECCOVID Given a query on COVID-19, retrieve documents that answer the query
BIOSSES Retrieve semantically similar text
SICK-R Retrieve semantically similar text
STS12 Retrieve semantically similar text
STS13 Retrieve semantically similar text
STS14 Retrieve semantically similar text
STS15 Retrieve semantically similar text
STS16 Retrieve semantically similar text
STS17 Retrieve semantically similar text
STS22 Retrieve semantically similar text
STSBenchmark Retrieve semantically similar text
SummEval Given a news summary, retrieve other semantically similar summaries
Table 14: MTEB instructions for evaluation of finetuned models.
Figure 11: Performance of the evaluated MTEB datasets for finetuning over the number of finetuning steps.
Model Average Clas. Clus. Pair Clas. Rera. Retr. STS Summ.
Classical (w/ instruct., mean) 62.96 76.26 42.68 86.31 57.58 53.75 81.53 30.19
Classical (w/ instruct., last) 63.98 76.57 45.78 86.37 56.71 54.87 82.03 31.02
Echo (w/ instruct., mean) 64.22 77.00 44.94 87.73 58.30 55.11 82.52 29.46
Echo (w/ instruct., last) 64.68 77.43 46.32 87.34 58.14 55.52 82.56 30.73
Classical (w/out instruct., mean) 62.19 75.23 41.79 85.24 56.31 53.24 80.97 30.64
Classical (w/out instruct., last) 62.37 75.01 42.70 85.69 56.64 53.29 80.92 30.91
Echo (w/out instruct., mean) 63.28 75.26 42.93 86.95 57.05 55.65 81.40 30.62
Echo (w/out instruct., last) 62.80 75.30 42.94 86.31 57.31 54.18 80.92 31.00
Classical (w/ instruct., mean, step 280) 63.19 76.18 42.99 85.44 57.63 53.96 82.53 29.94
Classical (w/ instruct., last, step 280) 63.87 76.54 46.22 86.70 57.79 53.73 82.22 30.13
Echo (w/ instruct., mean, step 280) 64.04 76.84 45.76 87.72 59.33 53.55 82.64 30.33
Echo (w/ instruct., last, step 280) 64.50 76.41 46.70 87.17 59.10 54.84 82.98 31.09
Table 15: Additional ablations for finetuning.
Dataset Repetition (last) Repetition (mean) Classical (last) Classical (mean) Bidirectional (last)
AmazonCounterfactualClassification 82.97 82.91 80.82 82.21 83.07
AmazonPolarityClassification 90.98 88.25 92.55 90.37 90.83
AmazonReviewsClassification 48.71 49.41 48.75 46.76 47.94
Banking77Classification 88.15 88.06 87.95 87.69 88.17
EmotionClassification 52.18 51.51 50.66 49.23 52.09
ImdbClassification 87.42 84.80 83.18 82.53 83.02
MassiveIntentClassification 79.67 79.70 78.60 79.15 78.93
MassiveScenarioClassification 82.82 82.74 81.71 81.46 81.80
MTOPDomainClassification 96.16 96.10 95.92 95.54 96.14
MTOPIntentClassification 85.75 85.87 85.96 85.86 85.98
ToxicConversationsClassification 71.91 72.21 71.19 72.21 71.46
TweetSentimentExtractionClassification 62.40 62.46 61.60 62.07 60.97
ArxivClusteringP2P 47.02 45.52 46.73 45.80 47.03
ArxivClusteringS2S 43.52 42.32 43.99 40.73 42.14
BiorxivClusteringP2P 35.53 35.24 36.50 35.42 36.21
BiorxivClusteringS2S 35.34 33.70 34.87 32.03 34.77
MedrxivClusteringP2P 30.27 29.68 30.67 29.74 31.06
MedrxivClusteringS2S 29.67 27.73 29.75 27.97 30.12
RedditClustering 61.77 59.12 61.17 54.79 62.50
RedditClusteringP2P 66.01 65.44 64.84 63.68 65.45
StackExchangeClustering 72.04 71.21 71.87 66.99 71.58
StackExchangeClusteringP2P 35.29 34.07 33.08 31.47 34.98
TwentyNewsgroupsClustering 53.04 50.29 50.07 40.91 49.53
SprintDuplicateQuestions 94.59 95.05 94.38 95.29 96.26
TwitterSemEval2015 79.93 80.73 77.18 75.98 80.80
TwitterURLCorpus 87.50 87.40 87.56 87.67 87.38
AskUbuntuDupQuestions 64.13 64.44 62.24 63.32 62.65
MindSmallReranking 32.92 32.11 32.68 32.52 32.53
SciDocsRR 83.68 84.15 81.60 83.01 82.36
StackOverflowDupQuestions 51.84 52.51 50.33 51.48 51.35
ArguAna 58.52 56.52 57.22 51.14 57.27
ClimateFEVER 34.56 37.07 31.10 30.31 32.73
CQADupstackRetrieval 46.91 46.48 45.11 43.30 46.52
DBPedia 46.83 48.19 45.18 46.80 46.76
FEVER 91.22 91.14 90.30 90.63 91.66
FiQA2018 54.51 54.11 50.31 48.94 53.06
HotpotQA 76.41 75.75 72.95 68.50 75.30
MSMARCO 43.25 43.11 42.31 41.49 43.38
NFCorpus 39.55 37.18 39.32 38.53 38.61
NQ 62.31 61.51 62.07 60.65 63.69
QuoraRetrieval 89.34 89.33 89.04 88.94 89.57
SCIDOCS 20.17 17.73 19.34 19.88 19.69
SciFact 73.99 73.57 74.22 75.39 75.83
Touche2020 18.52 18.92 24.46 19.44 15.79
TRECCOVID 76.66 76.02 80.17 82.30 74.50
BIOSSES 86.54 86.78 85.73 83.31 85.38
STS12 76.13 75.89 75.84 76.23 75.50
STS13 83.19 82.90 83.41 82.61 83.44
STS14 80.60 80.99 79.80 79.89 81.35
STS15 87.16 87.16 86.99 86.68 87.43
STS16 85.16 84.93 83.93 84.18 85.34
STS17 90.88 90.78 91.12 90.14 90.99
STS22 67.04 67.21 66.27 65.99 66.32
STSBenchmark 85.67 85.87 84.96 85.20 85.45
SICK-R 83.23 82.70 82.22 81.11 82.97
SummEval 30.73 29.46 31.02 30.19 29.32
Table 16: Results from all MTEB datasets for finetuning with Mistral-7B.

### v2

@@
 Abstract
-Recent approaches to improving the extraction of text embeddings from autoregressive large language models (LLMs) have largely focused on improvements to data, backbone pretrained language models, or improving task-differentiation via instructions.
-In this work, we address an architectural limitation of autoregressive models: token embeddings cannot contain information from tokens that appear later in the input.
-To address this limitation, we propose a simple approach, “echo embeddings,” in which we repeat the input twice in context and extract embeddings from the second occurrence.
-We show that echo embeddings of early tokens can encode information about later tokens, allowing us to maximally leverage high-quality LLMs for embeddings.
-On the MTEB leaderboard, echo embeddings improve over classical embeddings by over $9\%$ zero-shot and by around $0.7\%$ when fine-tuned.
-Echo embeddings with a Mistral-7B model achieve state-of-the-art compared to prior open source models that do not leverage synthetic fine-tuning data.
-[^1] Our code and pre-trained models are released at https://github.com/jakespringer/echo-embeddings.
-Repetition Improves Language Model Embeddings
-Jacob Mitchell Springer  Suhas Kotha Daniel Fried  Graham Neubig  Aditi Raghunathan Carnegie Mellon University {jspringe, suhask, dfried, gneubig, aditirag}@cs.cmu.edu
+Bidirectional models are considered essential for strong text embeddings.
+Recent approaches to adapt autoregressive language models (LMs) into strong text embedding models have largely had the requirement to modify the LM architecture to be bidirectional.
+We challenge this premise by introducing “echo embeddings” which converts autoregressive LMs into high quality text embedding models without changing the architecture or requiring fine-tuning.
+By repeating the input and extracting embeddings from the repeated tokens—which have access to all original tokens—echo embeddings improve over classical LM embeddings by over 5% in zero-shot settings.
+Our zero-shot embeddings nearly match those obtained by bidirectionally-converted LMs that undergo additional masked-language modeling training.
+Echo embeddings are also compatible with supervised fine-tuning, matching or outperforming bidirectionally-converted LMs in an apples-to-apples comparison, even with an identical compute budget during training and inference.
+Overall, repetition is a simple and effective strategy to circumvent the need for bidirectional attention in embedding models, paving the way towards a unified architecture for all NLP tasks.
 1 Introduction
+Figure 1: Overview of echo embeddings.
 Neural text embeddings have a crucial role in modern approaches to information retrieval (IR), semantic similarity estimation, classification, and clustering (Ni et al., 2021b; Muennighoff et al., 2022).
 For example, document retrieval often leverages low-dimensional embeddings for efficient lookup: when queries and documents are encoded as vectors where semantic relationships are described by similarity in some metric space, a query lookup can be reduced to an approximate nearest-neighbor search in embedding space (Johnson et al., 2019; Vanderkam et al., 2013).
-In the recent past, the dominant pretrained language model paradigm for neural embeddings have been masked language models with bidirectional attention (Ni et al., 2021a; Raffel et al., 2020; Izacard et al., 2021; Wang et al., 2022; Jiang et al., 2022; Su et al., 2022; Xiao et al., 2023a; Li et al., 2023).
-However, more recent literature (Ma et al., 2023; Wang et al., 2023) has begun to scale these algorithms to modern autoregressive language models such as LLaMA-2 and Mistral (Touvron et al., 2023; Jiang et al., 2023a).
-Developing approaches to construct embeddings from autoregressive language models is promising: for many tasks, these models are the highest quality models available (Srivastava et al., 2022).
-Figure 1: Conceptual overview of echo embeddings.
-In this paper, we address a striking failure mode of autoregressive language models.
-This failure arises from the fact that for autoregressive language models, contextualized token embeddings—the vector of last-hidden-layer activations at the position of a particular input token—do not contain information from tokens that appear later in the sentence due to the causal attention mask.
-We demonstrate that such embeddings can fail to appropriately determine similarity when the early tokens are superficially similar but become dissimilar in important ways when using key information from end of the input.
-We propose a strategy to overcome this limitation in autoregressive models through “echo embeddings.” With this approach, we repeat the inputs so that it appears twice in the context passed to the language model, and extract embeddings from the second occurrence.
-Repeating the input enables the contextualized token embeddings of the second occurrence of the passage to encode information from tokens that appear later in the passage by attending to their first occurrence in the passage.
-We show that echo embeddings do in fact allow embeddings of the early tokens to capture information about the later tokens.
-We then evaluate echo embeddings on the standard Massive Text Embedding Benchmark (MTEB) leaderboard [^2] The MTEB leaderboard can be found at https://huggingface.co/spaces/mteb/leaderboard.
-In the zero-shot setting, echo embeddings improve on classical embeddings by over $9\%$ and provide consistent gains across all the different tasks for a variety of language models and scale.
-We then perform an apples-to-apples comparison when fine-tuning embeddings from Mistral-7B and continue to see consistent gains of echo over classical across the various tasks (by $0.7\%$ on average).
-Strikingly, echo embeddings with the strong Mistral-7B language model allows us to achieve state-of-the-art embedding quality, enabling autoregressive language models to match prior open-source top performing models that otherwise leveraged MLMs with bidirectional attention.
-[^3] The contemporaneous work by Wang et al.
-(2023) achieves state-of-the-art accuracy on MTEB using classical embeddings from autoregressive language models, but through finetuning on high quality synthetic data, which we believe is largely orthogonal to our contribution.
-The approach of echo embeddings is conceptually well-motivated, extremely simple, and generally compatible with other innovations in extracting embeddings from autoregressive language models.
-As language models are likely to continue to improve over the coming years, echo embeddings can be a simple but powerful twist to classical embeddings that allow us to maximally leverage autoregressive language models.
+Until recently, the dominant pretrained language model paradigm for neural embeddings have been masked language models with bidirectional attention (Ni et al., 2021a; Raffel et al., 2020; Izacard et al., 2021; Wang et al., 2022; Jiang et al., 2022; Su et al., 2022; Xiao et al., 2023a; Li et al., 2023).
+Very recent literature has made progress generating high quality embeddings from autoregressive language models.
+However, these approaches either require large-scale synthetic data (Wang et al., 2023) or for the architecture of the base model to be “cast” to have bidirectional attention (BehnamGhader et al., 2024; Muennighoff et al., 2024).
+The requirement to change the architecture prevents the extraction of high quality embeddings without additional training.
+Further, this requirement is an obstacle to the development of truly unified models that can generate text and produce embeddings simultaneously without changing the architecture between encoding text for embeddings and generating tokens.
+Is bidirectional attention crucial for high quality embeddings?
+Intuitively, autoregressive models fail to capture information across the entire input due to the causal attention mask.
+We show this concretely in a toy setup in Section 3.
+In this work, we present a surprisingly effective way to overcome this limitation without requiring bidirectional attention or fine-tuning.
+Our method, “echo embeddings” is simple: we prompt the model with the input twice and extract embeddings from the second occurrence of the input.
+Even with causal attention, the embeddings of the second occurrence can access the entire input in the first occurrence.
+The second key ingredient is to prompt the language model to repeat the input (or, for example, to fix, rephrase, or otherwise reconstruct the input).
+This simple change enables the model to extract bidirectional information.
+As a result, echo embeddings offer strong performance in the entirely zero-shot setting—without any additional fine-tuning.
+In comparison with the classical method of extracting embeddings, the zero-shot performance of echo embeddings is around 5% higher.
+This nearly matches the performance of the recent more expensive method of casting to bidirectional attention followed with an addititional unsupervised MLM training (BehnamGhader et al., 2024).
+Even in comparison to the strong zero-shot baseline of PromptEOL (Jiang et al., 2023b)—a method in which the model is prompted to summarize the input in a single token—echo embeddings outperform by a significant margin, and is less sensitive to the exact choice of prompt.
+We emphasize that echo embeddings are simple, cheap, and flexible.
+The current best embeddings still require additional fine-tuning with a supervised objective.
+We find that echo embeddings slightly outperform the classical approach to extracting embeddings, even when the architecture is cast to be bidirectional, and even when echo embeddings is given an identical compute budget.
+We speculate that these gains are because echo embeddings better preserve and leverage large-scale pretraining because there are no architectural changes.
+This is in contrast to the bidirectional relaxation approach, which requires the model to adjust its weights to perform well with bidirectional attention.
+Echo embeddings come with a lot of benefits: simple to apply, works zero-shot without additional training, allows a unified model to be used for various tasks, and offers (slightly) better performance than casting to bidirectional attention offers.
+One apparent drawback is that repetition doubles the compute cost.
+To account for this, in all our experiments, we report “compute matched” results for both inference time and training time (when applicable).
+While the gains are now slightly smaller, we see that echo embeddings still perform well.
+In total, we believe that the echo embedding approach, by virtue of its simplicity and effectiveness, is an important step in bridging the gap between traditionally bidirectional embedding models and the current predominant paradigm of autoregressive language models.
 2 Preliminaries
 Our goal is to extract text embeddings that map a sentence $x$ to a vector $\phi(x)\in\mathbb{R}^{d}$ such that the semantic similarity between sentences is captured as similarity between their embeddings.
-In practice, we use the cosine similarity between embeddings to capture semantic similarity (detailed in Appendix B).
+In practice, we use the cosine similarity between embeddings as a similarity metric (detailed in Appendix C).
 Embeddings from language models.
@@
 We can pool the embeddings across all the tokens in different ways.
-In this work, we focus on two common strategies which have been considered by prior work (Reimers and Gurevych, 2019; Muennighoff, 2022; Zhang et al., 2023a; Wang et al., 2023).
+In this work, we focus on two common strategies which have been considered by prior work (Reimers & Gurevych, 2019; Muennighoff, 2022; Zhang et al., 2023a; Wang et al., 2023).
+Pooling.
 A mean token embedding over a set of indices $A$, refers to the mean contextualized token embeddings at indices in $A:\phi_{A}(x)\coloneqq\frac{1}{\left|A\right|}\sum_{t\in A}\phi_{t}(x)$.
@@
 We will refer to embeddings created in this way as “classical embeddings”.
-Additionally, one might first prompt the language model with an explanation of the task of interest followed by the sentence, and then pool the contextualized embeddings of the sentence tokens like before (Su et al., 2022).
-3 Echo Embeddings
-In this section, we first demonstrate a failure mode of classical embeddings, and motivate a new method that we call echo embeddings that addresses this failure.
-3.1 Classical Embeddings Miss Bidirectional Information.
-Sentence embeddings should aggregate information across the entire sentence.
-However, for autoregressive language models, the contextualized embedding at position $k$ $\phi_{k}(x)$ cannot encode information about tokens $x_{k+1},x_{k+2},\ldots$.
-Hence, the “meaning” encoded by the embeddings of tokens at the beginning of a sentence might inaccurately suggest they are similar (or dissimilar) to other tokens without considering the influence of tokens that come later.
-As a simple illustration, consider the following.
-$q$ $\colon{{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[She loves summer]}}}{{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [but dislikes the heat]}}}$
-$s^{-}$ $\colon{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[She loves summer]}}{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [for the warm evenings]}}$
-$s^{+}$ $\colon{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[Summer is her favorite]}}{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [but not the temp.]}}$
-Here, the contextualized embeddings of the first half of $s^{+}$ and $s^{-}$ are both similar to $q$ because they do not attend to the second half of the sentence.
-As a result, the similarity between $q$ and $s^{-}$ would be overestimated by any pooling strategy that uses information from the first half.
-We address last-token pooling at the end of this section.
-3.2 Echo Embeddings
-We propose a simple fix to mitigate the failure above: we present the input sentence twice to the language model and extract contextualized embeddings from the second occurrence of the sentence.
-In principle, the contextualized embeddings of the second occurrence can attend to the entire sentence presented in the first occurrence.
-Furthermore, in order to encourage the second occurrence to actually “encode” information about the first, we instruct the language model to perform a generic task that requires using this information, e.g., “rewrite” or “repeat.”
-Classical embeddings: Feed sentence $x$ to the language model and pool the contextualized embeddings of sentence $x$.
-Echo embeddings: Feed a prompt such as “Rewrite the sentence: $x$, rewritten sentence: xx” to the language model and pool the contextualized embeddings of the second occurence of $x$.
-Key to our method is passing the sentence twice to the model and pool embeddings exclusively from the second occurrence.
-[^4] We find that minor variations of the echo embeddings prompt (e.g.
-change “rewrite” to “repeat”) work equally well and we provide an example list in Appendix B Other tricks from classical embeddings such as prompting the model with the downstream task of interest can be applied to echo embeddings as well.
-Figure 2: (left) We take the echo embeddings of only AA for query $q$ and sentences $s^{-},s^{+}$ and plot the distribution of cosine similarities, showing that echo embeddings encode later information in earlier tokens.
-(right) We plot the accuracy of classical and echo embeddings when the sentences have similar beginnings (Structure 1) and dissimilar beginnings (Structure 2).
-3.3 Repetition Captures Bidirectional Info
-In the previous section, we argued that classical embeddings suffer from the issue that contextualized embeddings of early tokens can miss out on information from the later tokens.
-But can simply repetition via echo embeddings solve this issue?
-We aim to test this by extracting embeddings from Mistral-7B on a simple controlled synthetic setting.
-Given a query $q\colon[A,B]$ we construct sentence pairs $s^{+},s^{-}$ as follows. We make the first parts of each sentence identical to the query, but differ only in their second parts,
-$q\colon[A,B]$; $s^{+}\colon[A,B^{+}]$; $s^{-}\colon[A,B^{-}]$,
-where $B^{+}$ is $B$ but paraphrased and $B^{-}$ is semantically dissimilar to $B$. We query GPT-4 to generate examples of this structure. We describe the full procedure and prompts in Appendix B.
-With classical embeddings, the contextualized embeddings of $A$ parts of $s^{+},s^{-},q$ are identical by construction.
-To test whether echo embeddings can meaningfully distinguish $s^{+}$ and $s^{-}$ despite having identical initial tokens, we take the mean over just the $A$-portion of the echo embeddings and plot the cosine similarities $\operatorname{Sim}(q,s^{+})$ and $\operatorname{Sim}(q,s^{-})$ in Figure 2 (left).
+Additionally, one might first prompt the language model with an explanation of the task of interest followed by the sentence (Su et al., 2022).
+3 Echo Embeddings from LLMs
+Figure 2: A.
+Accuracy of classical embeddings (mean and last-token pooling) and echo embeddings on the entire toy dataset.
+B.
+Accuracies of classical and echo embeddings split by the different structures of the data (S1 and S2, see text).
+C.
+We take the echo embeddings of only the first parts $A$ of query $q=[A,B]$ and sentences $s^{-},s^{+}$ and plot the distribution of cosine similarities, showing that echo embeddings encode can later information in earlier tokens.
+In this section, we describe our general methodology to extract high quality embeddings from autoregressive language models without bidirectional attention.
+We start with a simple synthetic dataset that demonstrates why causal attention might inhibit embeddings from capturing information reliably across the entire context (Section 3.1).
+We then present our proposed method of Echo Embeddings in Section 3.2, and test that our simple and intuitive approach in fact addresses the failure mode identified in the synthetic setup without introducing bidirectional attention.
+3.1 Failure of classical embeddings with causal attention
+High-quality embeddings should effectively capture various text structures, including where similarity between inputs is strongly affected by either the early parts or the later parts of the inputs.
+We create a dataset with two “opposite” structures (S1 and S2) and show the classical embeddings with causal attention cannot capture both simultaneously.
+The dataset we construct consists of a query $q$ and a similar sentence $s^{+}$ and a dissimilar sentence $s^{-}$.
+We measure the performance of embedding $\phi(\cdot)$ by computing the fraction where $\operatorname{Sim}(q,s^{+})>\operatorname{Sim}(q,s^{-})$.
+Each query and sentence is a concatenation of two strings, allowing us to vary whether early (or later) parts are similar (or discriminatory).
+Notationally, $A,A^{+}$, $B,B^{+}$ are semantically similar, while $A,A^{-}$, $B,B^{-}$ semantically dissimilar.
+We query GPT-4 to generate these various realistic but synthetic strings (see Section C).
+S1 (Early discriminatory; late redundant).
+For a query $q=[A,B]$, in this structure, the early tokens provide information to discriminate between $s^{+},s^{-}$ while the later tokens are all similar.
+In other words, $s^{+}:[\mathbf{A^{+}},B^{+}]$ and $s^{-}:[\mathbf{A^{-}},B]$, where the discriminatory parts are bolded for convenience.
+S2 (Early redundant; late discriminatory).
+Analogously, we consider a second “opposite” structure where the early tokens are all similar but the later tokens can discriminate between $s^{+},s^{-}$.
+For query $q=[A,B]$, we have $s^{+}:[A^{+},\mathbf{B^{+}}]$ and $s^{-}:[A,\mathbf{B^{-}}]$, where the discriminatory parts are bolded for convenience.
+We show the results of zero-shot classical embeddings extracted from Mistral-7B-Instruct-v0.1 in Figure 2 (A) on a dataset that has a mixture of both structures.
+We see that neither mean nor last token pooling allows classical embeddings to achieve better-than-random performance.
+To get some insight into the failure, we plot the performance on each substructure separately in Figure 2 (B).
+Last-token pooling disproportionately weights the later parts of the sentence thereby missing discriminatory information in the early tokens in S1.
+In general, last-token pooling does not work well in a zero-shot manner on real data as well, as we see in the next section.
+Mean-token pooling is better suited to the zero-shot setting and can allow us to extract information from all parts of the input.
+This works well in structure S1.
+However, the causal attention produces a different issue: the contextual embedding at position $k$, $\phi_{k}(x)$ cannot encode information about the later tokens $x_{k+1},x_{k+2},\ldots$.
+This causes mean-token embeddings to exaggerate early token similarity and dilute discriminatory information in the later parts because of causal attention.
+This effect is especially exacerbated when the early tokens are similar, as in S2.
+In total, classical embeddings fail when the dataset contains inputs of both instances—where the discriminatory information could be in the early or later tokens.
+We show that this failure mode of classical embeddings can be overcome in a completely zero-shot manner using a simple method which we describe next.
+3.2 Echo Embeddings: A General Framework
+We saw in Section 3.1 that while mean-token pooling aggregates information across all tokens, the contextualized representations of early tokens are suboptimal because they cannot attend to the later tokens due to the causal attention.
+Replacing the causal with bidirectional attention is a natural solution, and common intuition seems to be that this is necessary.
+We propose a new general embedding framework—echo embeddings—that captures contextualized information from all tokens in the input with only causal attention:
+Echo embeddings method: Prompt the language model to act as an autoencoder, e.g., by asking the model to {repeat, rephrase, fix, etc.} the input; feed the sentence $x$ to the language model twice; pool the contextualized embeddings of the second occurrence of $x$.
+Ex: “Rewrite the sentence: $x$; rewritten sentence: $\bf{x}$”; embeddings are pooled from the bolded $\bf{x}$.
+The second occurrence can attend to everything, and hence extracting contextualized representations from the second occurrence could capture information from the entire input.
+Furthermore, in order to encourage the second occurrence to actually “encode” information about the first, we instruct the language model to perform a generic task that requires using this information, e.g., “rewrite” or “repeat”.
+The exact prompt, template, or task used to extract embeddings can be varied to suit the model and downstream task.
+We find empirically that the performance of echo embeddings are not particularly sensitive to the exact phrasing of the prompt (Section 4.2).
+We provide a list of example prompts in Appendix C.
+While echo embeddings is simple and intuitive, does it actually effectively capture bidirectional information?
+We apply echo embeddings with mean-token pooling on the synthetic dataset introduced in Section 3, and present the results in Figure 2.
+Echo embeddings with mean-token pooling works significantly better than classical embeddings and works well on both structures in the data—whether the discriminatory information is in the early or later tokens.
+In particular, let us focus on structure S2 where the discriminatory information appears in the later tokens.
+Echo embeddings with mean-token pooling work well, while classical embeddings did not because the early tokens could not attend to the later ones.
+This implies that echo embeddings can successfully address this failure and the representations of early tokens can capture information “bidirectionally” about the entire input without changing causal to bidirectional attention.
+To test this further, we measure whether echo embeddings can meaningfully distinguish $s^{+}$ and $s^{-}$ using just the information in the $A$ portions despite the discriminatory information solely appearing in the later $B$- parts.
+Recall that echo embeddings extract embeddings from the second occurrence.
+We plot the cosine similarities $\operatorname{Sim}(q,s^{+})$ and $\operatorname{Sim}(q,s^{-})$ in Figure 2 (C).
 We find that $\operatorname{Sim}(q,s^{+})$ is typically larger than $\operatorname{Sim}(q,s^{-})$.
-Since we are only pooling the echo embeddings of the A-portion, any distinction between $s^{+},s^{-}$ must come from the echo embeddings of $A$ capturing information from the later parts of the sentence.
-This showcases that current autoregressive language models can in fact allow early tokens to capture information from later tokens via echo embeddings.
-3.4 Classical vs. Echo on Synthetic Data
-In Section 3.3 we demonstrated that echo embeddings encode bidirectional information.
-However, is this sufficient to recover from the failure mode of classical embeddings?
-Further, where will we expect echo embeddings to improve over classical embeddings?
-Here, we compare echo and classical embeddings on synthetic data to answer both of these questions.
-Datasets.
-We sample datasets according to two structures depending on whether the discriminating information between $s^{+}$ and $s^{-}$ is in the first half (structure S1) or second half (structure S2) of the sentence.
-Using the structures below, we generate samples using GPT-4, as in the previous section (full details in the appendix):
-(S1) $q\colon[A,B]$, $s^{+}\colon[A^{+},B^{+}]$, $s^{-}\colon[A^{+},B^{-}]$
-(S2) $q\colon[A,B]$, $s^{+}\colon[A^{+},B^{+}]$, $s^{-}\colon[A^{-},B^{+}]$.
-We measure the accuracy of identifying which of two sentences $s^{+}$ and $s^{-}$ is closer to the query as measured by the cosine similarity in the embeddings.
-We compare classical vs echo embeddings when using mean pooling to aggregate embeddings extracted from Mistral-7B model.
-We use mean token embedding rather than last token because last token embeddings can be quite fragile in a zero-shot setting (Section 5.1).
-Results.
-We present results on the two different structures in Figure 2 (right).
-We see that classical embeddings struggle on Structure 1—when the distinguishing information is at the beginning, the embeddings corresponding to these early tokens exaggerate similarity between $q$ and $s^{-}$ because they do not encode the information provided by $B^{-}$.
-In contrast, echo embeddings are able to successfully determine the more similar sentence, presumably because embeddings of the $A$-portion now also encode information about the later parts (demonstrated in Section 3.3).
-As a control for other reasons echo embeddings outperform classical embeddings, we also compare these embeddings on structure two, where early tokens provide discriminative signal without needing the later context.
-As expected, both classical and echo embeddings achieve good performance in this setting.
-All in all, this analysis on synthetic data demonstrates that zero-shot classical embeddings do not encode information about later context in early token embeddings, but echo embeddings can do so.
-Does last-token pooling resolve the failure of classical embeddings?
-The embedding of the last token $\phi_{-1}(x)$ is, in principle, can encode information from the entire input.
-However, we posit that the last-token pooling strategy is highly brittle and can depend too strongly on the tokens near the end of the input.
-To verify this, we compare the accuracy of mean token pooling and last-token pooling for classical and echo embeddings in two settings.
-First, we evaluate on the original synthetic data of Structure 1 (Figure 3, left).
-Second, we evaluate on the same data, but where we append a uniformly randomly selected token to the end of each example (Figure 3, right).
-While last-token pooling has high accuracy on the original toy data (though still lower accuracy than echo embeddings), it fails to perform well on the noisy examples.
-Echo embeddings with mean token pooling, however, are robust to the noise.
-While this particular distribution of noise is artificial, it highlight that last-token pooling can be sensitive to noise in the last token.
-We verify in Section 5.1 that last-token embeddings perform poorly on real data.
-Thus, even if last-token embeddings address the inability of mean token classical embeddings to encode information from tokens that appear later in the sequence, they are not practical due to their sensitivity to noise.
-Does last-token pooling resolve the failure after finetuning?
-In practice, it is common to finetune embeddings on a sentence similarity objective.
-It is hard to delineate the degree to which this failure mode remains after finetuning.
-Nonetheless, we demonstrate in Section 5.2 that our method improves in the finetuning setting, even when using last-token pooling.
-Figure 3: We compare the accuracy of classical and echo embeddings, and mean and last-token pooling on sentences which have similar beginnings (Structure 1).
-We plot these accuracies on the original data (left), and the data in which a single uniformly randomly chosen token is appended to the end of each sentence (right).
-4 Methodology
-In Section 3, we explored how echo embeddings can improve over classical embeddings by addressing a fundamental failure mode.
-In this section, we describe the methodology by which we evaluate echo embeddings on large scale real datasets in both the zero-shot and finetuning settings.
-While finetuning is currently necessary to achieve state-of-the-art performance, zero-shot embeddings have the advantage that they do not require expensive finetuning on top of a pretrained language model.
-Zero-shot results can also more clearly show how different embedding strategies work on real datasets.
-4.1 Constructing Zero-shot Embeddings.
-We extract zero-shot embeddings via different strategies from three language models: Mistral-7B, LLaMA-2-7B, and LLaMA-2-13B.
+Since we are only pooling the echo embeddings of the $A$ portion, any distinction between $s^{+},s^{-}$ must come from the echo embeddings of $A$ capturing information from the later parts of the sentence.
+Summary.
+We construct a synthetic dataset that showcases a failure mode in extracting embeddings from causal language models in the classical manner.
+While ours is a simplistic construction, we demonstrate that the failure mode we outline does occur on real data as well in Appendix C and could explain the poor performance of classical embeddings in practice.
+However, echo embeddings, where we repeat the input and take embeddings from the second occurrence, can successfully address this failure mode.
+We now test echo embeddings on a variety of real-world settings.
+4 Experiments
+Table 1: Main results in the zero-shot setting: Apples-to-apples comparison of zero-shot echo and classical embeddings with mean pooling, and other baselines on MTEB (56 datasets).
+For the autoregressive embeddings in this table, we extract embeddings from the Mistral-7B-Instruct-v0.1 model.
+Bolding only zero-shot models.
+Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
+# of datasets 12 11 3 4 15 10 1 56
+Echo (Ours) 72.01 33.80 72.41 47.56 20.82 73.74 30.70 48.64
+Echo (compute matched) 71.63 33.51 72.31 47.43 22.85 73.64 31.02 49.02
+Classical 64.87 34.14 57.71 43.60 18.26 57.07 27.02 42.38
+PromptEOL 73.84 27.53 55.47 48.44 13.18 67.14 28.33 43.69
+BERT 61.47 29.48 56.33 43.44 10.44 52.98 29.82 37.87
+RoBERTa 62.63 29.05 56.95 41.92 8.62 55.24 28.64 37.86
+LLM2Vec (requires fine tuning) 72.51 40.06 70.95 50.43 19.74 71.90 27.84 49.43
+In this section, we describe how we implement and evaluate echo embeddings on real data.
+4.1 Massive Text Embedding Benchmark
+MTEB & MTEB-MINI.
+Our main evaluation dataset is the English-language subset of the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022).
+MTEB is a collection of 56 datasets that are grouped into different embedding tasks: classification, clustering, pair classification, reranking, retrieval, sentence similarity (STS), and summarization, with the goal of evaluating embeddings broadly.
+However, as the name suggests, MTEB is massive, and requires multiple days to evaluate a 7B-parameter model on 8 A100 GPUs.
+Thus, we adopt a smaller 28-dataset subset of MTEB that spans all categories except summarization, which we call MTEB-MINI.
+We provide a detailed description of MTEB and MTEB-MINI in Appendix A.1.
+4.2 Evaluation of Zero-shot Embeddings
+We begin with the zero-shot setting, where embeddings are extracted from a pre-trained model without additional training.
+Our goal is to show that without any additional training, echo embeddings outperform prior entirely-zero-shot methods, and can perform similarly to unsupervised approaches (which require additional fine-tuning with unsupervised data), such as LLM2Vec-unsupervised (BehnamGhader et al., 2024).
+While supervised fine-tuning is often used to improve embedders, zero-shot embeddings have the advantage that no further (potentially expensive) training is required.
+Table 2: Role of scale and base model in the zero-shot setting: Average MTEB score (56 datasets) for LLaMA-7B-2-Instruct and S-LLaMA-1.3B.
+Model Echo Classical PromptEOL
+S-LLaMA-1.3B 40.45 34.31 37.50
+LLaMA-7B 45.84 40.29 42.84
+Table 3: Casting to bidirectional attention: We evaluate MTEB-MINI (28 datasets) and compare classical with causal attention (Cl.+Uni) and with bidirectional attention (Cl.+Bi).
+Model Echo Cl.+Uni Cl.+Bi
+LLaMA-2-7B 56.99 47.27 43.03
+S-LLaMA-1.3B 49.55 40.14 35.77
+Mistral-7B 59.78 49.03 58.24
+Constructing zero-shot echo embeddings.
+Following the echo embeddings method described in Section 3, to construct echo embeddings we need to pick a prompt, a pooling strategy, and a language model.
+We evaluate the impact of these choices on the performance of the embeddings on the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022).
+Choosing a prompt: As specified in Section 3, we must prompt the language model to act as an autoencoder.
+Sclar et al.
+(2023) argues that on many tasks, the exact formatting and wording of the prompt can strongly influence performance.
+Thus, we investigate: does the choice of prompt affect embeddings quality?
+As we shall discuss later, we find that, the choice of prompt does not substantially affect the performance of echo embeddings (Section 4.3).
+This implies that prompt tuning is unnecessary, and any reasonable prompt will perform well.
+For our main results, we adopt the prompt “Rewrite the following paragraph: $S$.
+The rewritten paragraph: SS” for echo embeddings and “Write a paragraph: SS” for classical embeddings.
+Table 4: Role of pooling in the zero-shot setting: Average MTEB Score (56 datasets) for mean and last token pooling.
+Mean Last-token
+Echo 48.64 31.55
+Classical 42.38 31.94
+Choosing a pooling strategy: We evaluate both mean and last-token pooling for both classical and echo embeddings 4.
+Consistent with our toy-data analysis in Section 3, we find that mean token pooling is required for good embeddings and thus use this for all further zero-shot experiments (Table 4).
+Choosing a language model: We evaluate the performance of echo embeddings on three language models: Mistral-7B-Instruct-v0.1 (Jiang et al., 2023a), LLaMA-2-7B-Instruct (Touvron et al., 2023a), and S-LLaMA-1.3B (Xia et al., 2023) (sometimes abbreviated).
 We select the instruction-finetuned model for each of them.
 Refer to Appendix A.2 for additional information on the base models.
-Recent literature suggests that the performance of language models on zero-shot tasks can be highly variable depending on the exact wording and template of the prompts (Sclar et al., 2023).
-Thus, for each of the embedding strategies we consider, we perform prompt randomization where we sample prompts by randomizing the exact wording, punctuation, and capitalization of the prompt.
-We describe the sampling process and the exact prompts that we use in Appendix C.
-Baselines.
-We compare our proposed echo embeddings (Section 3.2) to classical embeddings and two additional baselines:
-  • Last-token embeddings: We mentioned in the Section 3 that last-token embeddings tend to underperform in comparsion to mean token embeddings, and thus we compare on real data.
-  • Summarization: We also compare zero-shot embeddings obtained via the strategy proposed by Jiang et al.
-(2023b).
-Here, they instruct the model to summarize the input in a single word and then take the last token embedding $\phi_{-1}(x)$ as the pooled embedding of the sentence.
-4.2 Constructing Finetuned Embeddings.
-We adopt the conventional sentence embedding training setup (Reimers and Gurevych, 2019) where we train with a contrastive learning objective to encourage the embeddings of similar text to be close.
-We extract embeddings in a slightly different fashion compared to the zero-shot setting above in order to keep the finetuning methodology as similar as possible to the existing literature.
-Extracting embeddings.
-The training and evaluation data is separated into two categories: symmetric data, in which sentences are drawn from a single distribution (such as for sentence similarity), and asymmetric, in which the data consists of both queries and documents (such as for retrieval).
-We adopt a separate prompt for symmetric inputs and queries, and for documents.
-We construct classical embeddings by encoding text $S$ using the following prompts:
-Queries & Symm. Documents
-Instruct: {instruction}
+Figure 3: Role of prompt in the zero-shot setting: Variance of zero-shot performance across different prompts and formats on MTEB-MINI (28 datasets).
+Refer to Appendex D for details on the exact prompts and formats.
+Compute-matching echo embeddings.
+Naively, echo embeddings would require twice the compute of classical embeddings.
+To account for this, we also consider compute-matched echo embeddings in which we halve the length of the input.
+Alternative approaches to constructing embeddings from autoregressive models.
+In addition to our results, we compare to two recently proposed methods to convert autoregressive models into embedders without supervised fine-tuning: PromptEOL (zero-shot) and the unsupervised variant of LLM2Vec (requires additional fine-tuning).
+PromptEOL prompts the model to summarize the input in a single token, and extracts embeddings from the predicted token (Jiang et al., 2023b) (see Appendix D for additional details).
+The unsupervised variant of LLM2Vec (LLM2Vec-unsupervised) incorporates an additional masked-language-modeling fine-tuning step.
+Importantly, LLM2Vec is not zero-shot because it involves this extra fine-tuning step.
+4.3 Zero-shot Results
+Echo embeddings yield strong embeddings, and without additional compute.
+We present the main zero-shot results in which we evaluate echo embeddings on MTEB in Table 1.
+We find that echo embeddings successfully converts causal language models into strong encoders.
+For Mistral-7B-Instruct, we find, surprisingly, that compute-matched echo embeddings, achieves slightly higher performance than full-compute echo embeddings.
+The difference is small enough that it is unclear to what degree we would observe this improvement in other settings, but suggests that echo embeddings can be adapted to outperform classical embeddings without any additional compute.
+We extend this analysis of compute in Section 4.6.
+We also evaluate echo embeddings over multiple models and scales in Table 3 and find consistent results.
+Comparison to PromptEOL.
+We find that echo embeddings outperform PromptEOL on average across the tested MTEB datasets by 5%.
+In addition to this improvement, echo embeddings is additionally much less sensitive to the exact wording and formatting of the prompt (Figure 3).
+Comparison to prior encoder models.
+We evaluate the zero-shot MTEB performance of existing bidirectional (encoder) models, BERT-large and RoBERTa-large, which are often used as the backbone of fine-tuned embedders.
+We find that echo embeddings outperform both BERT-large and RoBERTa-large in the zero-shot setting.
+We suspect that the performance improvement is due to Mistral-7B being significantly more powerful than BERT and RoBERTa.
+This highlights the importance of our method: by using echo embeddings, we can leverage the full power of autoregressive language models, which are generally dominant over bidirectional models on other NLP tasks.
+Comparison to LLM2Vec.
+We find that entirely zero-shot echo embeddings can perform similarly to LLM2Vec-unsupervised, despite the fact that LLM2Vec-unsupervised is not a zero-shot method, and requires additional fine-tuning on top of the pre-trained base model.
+Echo embeddings thus offers a simple alternative to LLM2Vec that does not require additional training.
+Relaxing to bidirectional attention.
+We find that changing the attention mask from causal to bidirectional often substantially underperforms the causal architecture, but not for Mistral-7B (Table 3).
+Regardless, echo embeddings consistently outperforms this baseline.
+A priori, we might expect that, without further fine-tuning, such a modification to the architecture would harm performance, as we would expect that the model has not been trained to accommodate the additional attention.
+For LLaMA-2-7B and S-LLaMA-1.3B, this is the case.
+However, for Mistral-7B, casting to bidirectional attention does not harm performance and instead yields performance similar to echo embeddings.
+We suspect that this may arise from non-standard pre-training methodology, as discussed by BehnamGhader et al.
+(2024).
+Ablations. We perform multiple ablations to evaluate the role of each component:
+  1. 1.
+    We test the role of pooling strategy in Table 4.
+Consistent with our analysis in Section 3, we find that mean pooling outperforms last-token pooling by a wide margin for both echo and classical embeddings.
+This suggests that mean pooling is required for echo embeddings in the zero-shot setting.
+  2. 2.
+    We evaluate the role of the exact wording and formatting of the prompt in Figure 3.
+We find that the performance of echo embeddings has low variance with respect to the exact choice of prompt.
+This implies that echo embeddings does not require prompt tuning to perform well.
+Further, we find that all tested echo embeddings prompts outperform every tested classical embeddings prompt.
+4.4 Evaluation of Fine-tuned Embeddings
+Echo embeddings also perform well in the fine-tuning setting.
+We will show that echo embeddings can outperform classical embeddings, even when the attention is converted to be bidirectional.
+Further, echo embeddings can match the performance of other methods that require much more complicated fine-tuning methodology, with identical fine-tuning and inference-time compute, and without modifying the architecture.
+Table 5: Main fine-tuning results: We report MTEB scores for echo embeddings with full compute and compute-matched and classical embeddings with a causal and with a casted bidirectional architecture.
+For echo embeddings, we use Mistral-7B as the backbone model, and last-token pooling.
+For the compute matched echo embeddings, FT+IT refers to both fine-tuning and inference-time compute matching.
+Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
+# of datasets 12 11 3 4 15 10 1 56
+Echo (ours) 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
+Echo (FT+IT compute matched) 77.39 46.27 87.49 58.08 55.07 83.18 30.73 64.66
+Classical 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
+Classical + bidirectional attn. 76.70 45.94 88.15 57.23 54.96 82.42 29.32 64.23
+Constructing fine-tuned echo embeddings.
+To construct fine-tuned embeddings, we must again choose a prompt, a pooling strategy, and a language model.
+In addition, we need to choose a contrastive learning objective and a collection of training datasets.
+Choosing a prompt: We adopt a minimal prompt that differs for queries and documents.
+Prior work has shown that the addition of an instruction can improve the quality of the embeddings (Wang et al., 2023), and thus we include an instruction in the prompt.
+The prompts are: Echo (queries) Echo (docs) Classical (queries) Classical (docs) Instruct: {instruction}
+$\text{Query: }S$ $\text{Query again: }S^{\prime}$
+$\text{Document: }S$ $\text{Document again: }S^{\prime}$ Instruct: {instruction}
 $\text{Query: }S$
 $\text{Document: }S$
-For echo embeddings, we use the prompts, where $S$ represents the input and $S^{\prime}=S$:
-Queries & Symm. Documents
-Instruct: {instruction}
-$\text{Query: }S$ $\text{Query again: }S^{\prime}$
-$\text{Document: }S$ $\text{Document again: }S^{\prime}$
-In this case, {instruction} refers to the task instruction, which specifies a description of the task that the embedding will be used for.
-We adopt the instructions from Wang et al.
-(2023), and provide a list of the instructions in Appendix D.
-We append an end-of-sentence token to the end of each input, and we allow the input embedding of this token to be trainable.
-Datasets.
-We train on a collection of publicly available datasets that encompass both symmetric and asymmetric data that are standard training datasets in the embedding literature.
-We list and describe each of the datasets in Appendix D.
-Optimization.
-To finetune the model, we optimize the SimCSE loss with in-batch and mined hard negatives.
-Since this is standard, we defer discussion of this to Appendix D.
+We provide a list of the instructions in Appendix F.
+Choosing a pooling strategy: We evaluate both mean and last-token pooling for both classical and echo embeddings.
+In practice, we find that last-token pooling performs slightly better than mean pooling in this setting (see Section 4.5).
+Choosing a language model: We train and evaluate the performance of echo embeddings on Mistral-7B-Instruct-v0.1 and S-LLaMA-1.3B (Jiang et al., 2023a; Xia et al., 2023).
+We choose Mistral-7B as the best-performing model in the zero-shot setting, and S-LLaMA-1.3B to test the broad applicability of our method across model class and scale.
+Table 6: Role of base model in the fine-tuning setting: We compare echo and classical embeddings for S-LLaMA-1.3B.
+FT+IT refers to the setting in which echo embeddings are compute-matched at train and inference time.
+Methods MTEB Score
+Echo 62.01
+Echo (FT+IT compute matched) 61.65
+Classical 61.26
+Table 7: Comparisons to recent baselines: MTEB score for models trained on similar data.
+Echo embeddings matches the performance of models trained on similar data, but without requiring bidirectional attention.
+FT+IT refers to the setting in which echo embeddings are compute-matched at train and inference time.
+Method MTEB Score
+Echo (FT+IT compute matched) 64.66
+LLM2Vec-supervised 64.80
+GritLM-7B (public data) 64.70
+Training datasets and optimization: We train on a collection of publicly available datasets that are standard training datasets in the embedding literature.
+We list and describe each of the datasets in Appendix F.
+To fine-tune the model, we optimize the SimCSE loss with in-batch and mined hard negatives.
+Since this is standard, we defer discussion of this to Appendix F.
 Each batch is constructed by sampling a dataset from our set of training dataset, and then collecting examples from only this dataset.
@@
 We choose $\tau=1/50$ and a learning rate of $8\times 10^{-4}$.
-We use the Mistral-7B instruction-tuned model as a backbone (Jiang et al., 2023a).
-Our choices aim to be consistent with prior literature (Wang et al., 2023; Su et al., 2022; Zhang et al., 2023a).
-4.3 Massive Text Embedding Benchmark
-For evaluation, we use the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022).
-For this paper, we focus on the English-language subset of the benchmark.
-MTEB is composed of a collection of 56 datasets that are grouped into different embedding tasks: classification, clustering, pair classification, reranking, retrieval, sentence similarity (STS), and summarization.
-The goal is to construct general purpose embeddings that are useful for solving each of the tasks.
-More information about MTEB is specified in Appendix A.1.
-For the fine-tuning setting, we evaluate on the entire English-language subset. In the zero-shot setting, for convenience, we only evaluate on a subset of MTEB. We describe this subset in Appendix D.
-5 Experiments
-Strategy Model Pool  Clas. P. Cls. Clus. Retr. STS Rera. Average
-Main results:
-Echo (ours) Mistral 7B Mean   64.06 75.26 27.02 23.61 72.40 60.00 55.07
-Classical Mistral 7B Mean   58.21 73.87 23.85 20.35 56.97 54.44 45.88
-Prior work:
-Summarization Mistral 7B Last 66.01 81.82 26.48 19.13 70.13 66.24 54.96
-Ablations:
-Echo Mistral 7B Last   63.11 57.93 12.82 2.97 39.14 47.35 36.60
-Classical Mistral 7B Last   58.23 46.64 13.51 2.60 33.97 46.51 32.52
-Echo LLaMA 7B Mean   61.64 66.29 25.11 16.12 66.18 56.35 50.26
-Classical LLaMA 7B Mean   56.61 68.46 23.22 18.63 56.49 53.26 44.65
-Echo LLaMA 13B Mean   64.65 74.57 25.72 26.58 72.20 62.68 55.60
-Classical LLaMA 13B Mean   58.50 65.06 24.22 18.92 57.47 56.38 45.15
-Table 1: Zero-shot scores on MTEB tasks for Mistral-7B.
-We use a retrieval validation set (FiQA2018) to select the best prompt.
-Refer to Appendix C for the scores with alternative validation sets.
-Top: Comparison of echo embeddings to classical embeddings.
-Center: Summarization approach to constructing embeddings (Jiang et al., 2023b).
-Bottom: Ablations, including last-token pooling and LLaMA-2-{7B, 13B}.
-5.1 Evaluation of Zero-shot Embeddings
-We compare the performance of classical, echo, and summarization embeddings on MTEB tasks (Table 1).
-We validate using a retrieval dataset from MTEB (FiQA2018) as a validation set as described in Section 4.3.
-We report the scores using alternative validation sets in Appendix C.
-Echo embeddings outperform classical embeddings zero-shot.
-We see that echo embeddings outperform classical embeddings by a large margin: on average, by nearly 10 points for Mistral-7B.
-Further, this performance increase is consistent across every MTEB category, across models (LLaMA-2 vs Mistral), and across scale (7B vs 13B).
-This demonstrates that echo embeddings can significantly improve the performance of embeddings on real data, suggesting that the failure mode of classical embeddings that we describe in Section 3 can affect performance on real data.
-Qualitative comparison of classical and echo embeddings.
-In Section 3, we demonstrate that classical embeddings overestimate the similarity between examples which are superficially similar based on tokens that appear early in the sequence.
-To build intuition that this applies to realistic data, we present the sentence pair from STSBenchmark, a sentence similarity task from MTEB, in which echo embeddings reduce error the most:
-[…]
-[…]
-which has a ground-truth score of 0 (out of $5$) similarity. The sentence pair for which echo embeddings reduces error the least is:
-[…]
-[…]
-which has a ground-truth similarity of $2$ (out of $5$). We provide more examples in the Appendix Table 7.
-For this example, notice, that the sentence pair $(x_{1},x_{2})$ on which echo embeddings improve error the most has exactly the property we identify as a failure mode for classical embeddings: the sentence is superficially similar for the first few tokens.
-On the other hand $(y_{1},y_{2})$ does not have this property.
-Quantitative evaluation of the failure mode.
-The above example builds intuition that, even on real data, classical embeddings fail to properly estimate similarity on examples which are superficially similar in the early tokens.
-We quantitatively measure the degree to which classical and echo embeddings fail on sentences which are similar for early tokens, and for sentences which are not.
-We find that classical embeddings systematically fail on examples which exhibit this structure, while echo embeddings do not.
-For convenience, we defer the discussion of these experiments and the results to Appendix C.1.
-Last-token vs mean token pooling.
-We find that last-token embeddings are substantially worse than mean token embeddings in the zero-shot setting, despite the fact that in principle, the last token in the sequence can encode information from all other tokens.
-In practice, it is clear that the last token does not encode sufficient information to achieve strong performance on MTEB in the zero shot setting.
-Echo embeddings vs summarization.
-We find that the average performance across the tested MTEB datasets is similar between echo and classical embeddings.
-Summarization does encourage the last token in the sequence to encode information about the entire sentence.
-We find that summarization is much more sensitive to the exact prompt while echo embeddings are robust to such minor variations (see Figure 5 in Appendix C).
-We suspect that echo embeddings are more robust as a result of more directly trying to encode bidirectional information into the embeddings.
-5.2 Evaluation of Finetuned Embeddings
-Strategy Model Pool   Clas. Clus. P. Cls. Rera. Retr. STS Average
-Main results:
-Echo (ours) Mistral 7B Last 77.43 46.32 87.34 58.14 55.52 82.56 64.68
-Classical Mistral 7B Last  76.57 45.78 86.37 56.71 54.87 82.03 63.98
-Prior work:
-UAE-Large-V1 (MLM)  75.58 46.73 87.25 59.88 54.66 84.54 64.64
-multilingual-e5-large (MLM)  77.56 47.10 86.19 58.58 52.47 84.78 64.41
-bge-large-en-v1.5 (MLM)  75.97 46.08 87.12 60.03 54.29 83.11 64.23
-udever-bloom-7b (autoregr.)  72.13 40.81 85.4 55.91 49.34 83.01 60.63
-sgpt-5.8b (autoregr.)  68.13 40.34 82.00 56.56 50.25 78.10 58.93
-e5-mistral-7b555e5-mistral-7b was recently released and leverages high quality synthetic data to achieve strong performance which is not publicly released.
-We report their performance, but we do not explicitly compare to them (Wang et al., 2023).
-(autoregr.)  78.47 50.26 88.34 60.21 56.89 84.63 66.63
-Ablations:
-Echo Mistral 7B Mean  77.00 44.94 87.73 58.30 55.11 82.52 64.22
-Classical Mistral 7B Mean  76.26 42.68 86.31 57.58 53.75 81.53 62.96
-Classical Mistral 7B-bidir. Last  76.70 45.94 88.15 57.23 54.96 82.42 64.23
-Table 2: Finetuning scores on MTEB tasks.
-Top: Apples-to-apples comparison of echo embeddings and classical embeddings in which we use echo embeddings and classical embeddings with last-token pooling, with the same training setup.
-Center: Performance of recent open source embedding models, annotated by base model type, masked-language model or autoregressive.
-Bottom: Ablations for finetuning: using mean token embeddings (first two lines) and using a bidirectional architecture (last line).
-Different embeddings on the MTEB leaderboard are often fine-tuned on different datasets.
-In order to perform an apples-to-apples comparison between embedding strategies, we fine-tune both echo and classical embeddings on the exact same datasets (described in Section 4.2).
-We report the results in Table 2.
-This table also includes a comparison to prior state-of-the-art methods using masked language models (MLM) and autoregressive language models.
-Further, we evaluate a number of ablations to determine the role of pooling strategy and architecture.
-Echo embeddings outperform classical embeddings after finetuning.
-We observe that echo embeddings consistently outperform classical embeddings on each category even after finetuning.
-Hence, the fundamental gap we find between classical and echo embeddings in Section 3 and in our zero-shot experiments persist after fine-tuning.
-Comparison to prior state-of-the-art models.
-We present comparisons to both prior MLM-based embeddings and prior autoregressive-language-model embeddings, listing the open-source models from the MTEB leaderboard.
-It is striking that MLMs vastly outperformed autoregressive models until recently.
-Our classical embeddings outperform the previous-best autoregressive language model.
-This is a result of using the strongest public 7B parameter language model (Mistral) and more fine-tuning data.
-However, despite these choices, classical embeddings do not outperform prior MLM-based approaches, perhaps because MLMs encode bidirectional context unlike classical embeddings from autoregressive models.
-Interestingly, echo embeddings allow us to close the gap to achieve state-of-the-art (on average) with an autoregressive model compared to prior open-sourced models on the leaderboard that used MLMs.
-A recent exception is the concurrent work by Wang et al.
-(2023) that use synthetic data to improve classical embeddings extracted from Mistral-7B.
-Their synthetic data is not publicly available, but the apples-to-apples comparison between classical and echo embeddings we performed suggests that echo embeddings could provide gains further gains over the numbers reported in (Wang et al., 2023) when fine-tuning with synthetic data.
-Why doesn’t last-token pooling close the gap?
-Since classical last-token embeddings can attend to every other token, they do not necessarily suffer from the failure mode that we highlighted in Section 3.
-Last token does not reliably capture relevant information in a zero-shot setting, but this could have been bridged via fine-tuning.
-It is thus surprising that, even after finetuning last-token embeddings that could (in principle) encode any embedding function, echo embeddings outperform classical embeddings.
-We identify two hypotheses that may explain this performance gap: (1) While last-token embeddings can attend to every token, the intermediate representations of earlier tokens cannot.
-If last-token pooling derives information from the internal representations of earlier tokens, by attending to these representations, last-token classical embeddings may still suffer from the failure mode of the earlier tokens.
-(2) If the post-finetuning performance benefits from the model initialization point, last-token classical embeddings may suffer: in Section 5.1 we show that last-token embeddings achieve poor zero-shot performance.
-We leave it to future work to explore these hypotheses.
-We do, however, observe that the gap between last- and mean token echo embeddings is smaller than the gap between last- and mean token classical embeddings, suggesting that echo embeddings can especially improve the quality of mean token embeddings.
-Can we relax autoregressive language models to a bidirectional architecture and fine-tune?
-To test the role of architecture, we finetune Mistral-7B on the same setup described in Section 4.2 but modified the architecture so as to remove the causal attention mask.
-While the initial weights are identical to Mistral-7B, this new model has bidirectional attention.
-We observe that the performance of bidirectional classical embeddings are better than our standard (causal) classical embeddings, but worse than echo embeddings.
-This suggests that the architecture alone is not sufficient to improve performance.
+Compute-matching echo embeddings.
+Echo embeddings require twice the compute of classical embeddings.
+Similar to the zero-shot setting, we reduce the inference-time cost of echo embeddings by halving the length of the input.
+However, in the training setting, we additionally must match the train compute.
+We opt to train for half as many steps as classical embeddings, as this is the most straightforward way to match the compute.
+The compute-matched echo embeddings in Table 5 refer to the setting where we match compute for both training and inference.
+Classical embeddings with bidirectional attention.
+In addition to our comparison to classical embeddings with causal attention, we train a model for which we replace the causal attention with bidirectional attention.
+4.5 Fine-tuning Results
+Echo embeddings yield high quality embeddings in the fine-tuning setting.
+We present the main fine-tuning results in Table 5.
+Similar to the zero-shot setting, compute-matched echo embeddings outperform classical embeddings with both architectures: causal and bidirectional.
+When additional compute is available, echo embeddings improve performance further.
+This suggests that the fundamental gap between classical and echo embeddings that we identified in Section 3 persists even after fine-tuning.
+We find consistent results when evaluating across model classes and scales (Table 7).
+Figure 4: Zero-shot setting: MTEB-mini score of zero-shot echo embeddings and classical embeddings with bidirectional attention with fixed compute (token) budget.
+We use the zero-shot Mistral 7B from Table 1.
+The standard token budget for MTEB is 512.
+Comparison to LLM2Vec.
+We compare to the supervised variant of LLM2Vec.
+The LLM2Vec method relaxes the attention of the base autoregressive model to bidirectional and requires two fine-tuning stages: (1) fine-tuning with an unsupervised masked-language-modeling objective (2) supervised fine-tuning for embeddings.
+Despite the complexity of the LLM2Vec method and additional fine-tuning, echo embeddings nonetheless approximately matches the performance with identical fine-tuning and inference-time compute (Table 7).
+Echo embeddings thus offers itself as a much simpler method that does not require modifying the architecture nor additional fine-tuning stages, with nearly identical performance.
+Comparison to GritLM.
+We also compare to GritLM (Muennighoff et al., 2024) in Table 7.
+This approach, which simultaneously trains an embedder and a generative model, finds the bidirectional attention relaxation to be crucial.
+We find that compute-matched echo embeddings perform similarly the variant of GritLM that has been trained on similar data as ours (public data), even though echo embeddings use a causal architecture.
+The role of pooling.
+We investigate the role of pooling in Appendix F.
+In the fine-tuning setting, the choice of pooling strategy has a much weaker effect than in the zero-shot setting.
+Consistent with prior work (Wang et al., 2023; Zhang et al., 2023a), we find that last-token pooling, with the addition of a trainable end-of-sentence token embedding, slightly outperforms mean pooling.
+Figure 5: Fine-tuning setting: MTEB-mini score of fine-tuned echo embeddings and classical embeddings with bidirectional attention with fixed compute (token) budget.
+We use the trained Mistral-7B models from Table 5.
+The standard token budget for MTEB is 512.
+4.6 Addressing the Inference-Time Cost of Echo Embeddings
+One potential drawback of echo embeddings is the computational cost of the method at both training and inference time.
+Naively, echo embeddings requires approximately twice the compute of classical embeddings, as the input is repeated.
+We have so far seen that in an entirely compute-matched training and inference time setting, echo embeddings outperforms classical embeddings (Table 5).
+In this section, we expand this analysis to consider how echo embeddings perform under different inference-time compute budgets.
+We explore the role of training under different budgets in Appendix F.
+4.7 The role of compute
+The role of inference-time compute in the zero-shot setting.
+In Section 4.2, we demonstrated that by halving the number of input tokens that are encoded by echo embeddings, thus matching compute with the classical approach to emebddings, we achieve stronger performance than classical embeddings.
+Here, we investigate the role of compute in further detail by plotting the performance of echo embeddings as a function of the total number of tokens encoded (Figure 4).
+We find that when the budget is greater than approximately 64 tokens, echo embeddings outperforms Mistral-7B classical embeddings with bidirectional attention.
+For very small computational budgets (fewer than 64 tokens), we find that Mistral-7B classical embeddings with bidirectional outperform echo embeddings.
+Since the standard budget for MTEB is 512 tokens, echo embeddings likely offer stronger performance in most settings.
+Over all compute budgets, echo embeddings outperform causal attention classical embeddings.
+The role of inference-time compute in the fine-tuning setting.
+The fine-tuning setting exhibits a similar trend: for very small computational budgets, echo embeddings perform worse than classical embeddings, but as the budget increases echo embeddings outperform classical embeddings (Figure 5).
+In the fine-tuning setting, echo embeddings outperform classical embeddings for budgets greater than 128 tokens, which is still much smaller than the standard MTEB budget (512 tokens).
 5 Related Work
@@
 Early approaches involved computing embeddings for individual words (Hinton, 1984; Rumelhart et al., 1986; Elman, 1990; Mikolov et al., 2013; Pennington et al., 2014).
-Later work aims to compute dense vectors representing the semantics of entire sequences by combining or composing word vectors (Le and Mikolov, 2014; Iyyer et al., 2015; Kiros et al., 2015; Socher et al., 2011; Tai et al., 2015; Wang et al., 2016; Wieting et al., 2015).
-Khattab and Zaharia (2020) propose to use late interaction between document and query vectors to improve retrieval performance.
-Reimers and Gurevych (2019) propose S-BERT which takes a pretrained BERT (Devlin et al., 2018) and trains with a triplet loss on anchor sentences, semantically similar positive examples, and semantically dissimilar negative examples.
-More recent approaches typically adopt this approach with different pretrained models and a contrastive objective such as InfoNCE (Oord et al., 2018) or SimCSE (Gao et al., 2021b).
-Ni et al.
-(2021a) with Ni et al.
-(2021b) extend this approach to the T5 architecture (Raffel et al., 2020).
-Multiple papers use an additional unsupervised contrastive objective (Izacard et al., 2021; Wang et al., 2022).
+Later work aims to compute dense vectors representing the semantics of entire sequences by combining or composing word vectors (Le & Mikolov, 2014; Iyyer et al., 2015; Kiros et al., 2015; Socher et al., 2011; Tai et al., 2015; Wang et al., 2016; Wieting et al., 2015).
+Khattab & Zaharia (2020) propose to use late interaction between document and query vectors to improve retrieval performance.
+Reimers & Gurevych (2019) propose S-BERT which takes a pretrained BERT (Devlin et al., 2018) and trains with a triplet loss.
+More recent approaches typically adopt this approach to different pretrained models (Raffel et al., 2020; Ni et al., 2021b; Raffel et al., 2020) and contrastive objectives (Oord et al., 2018; Gao et al., 2021b).
 Other papers propose including prompts to improve task-specific embedding performance (Jiang et al., 2022; Su et al., 2022).
 Some work combines multiple of these training objectives and approaches (Xiao et al., 2023a; Li et al., 2023).
-Notably, except for the most recent approaches, nearly all embeddings were based upon bidirectional architectures that were often pretrained with a masked-language modeling objective.
+Most recent sentence embeddings research has focused on improving finetuning.
+Reimers & Gurevych (2019) demonstrates that without finetuning, BERT has low-quality embeddings.
+To our knowledge, Jiang et al.
+(2023b) is the only paper that constructs zero-shot embeddings for autoregressive language models.
 Next-token language modeling for embeddings.
 A series of papers aim to construct high quality embeddings from autoregressive large language models.
-Multiple papers apply the fine-tuning approach of S-BERT to language models but using a trained GPT (Radford et al., 2018) as the backbone architecture (Muennighoff, 2022; Zhang et al., 2023a).
+Most closely related, BehnamGhader et al.
+(2024) proposed LLM2Vec which adds an additional unsupervised training step on a masked-language-modeling objective.Multiple papers apply the fine-tuning approach of S-BERT to language models but using a trained GPT (Radford et al., 2018) as the backbone architecture (Muennighoff, 2022; Zhang et al., 2023a).
 Ma et al.
-(2023) adopts this approach but for LLaMA-2 (Touvron et al., 2023).
+(2023) adopts this approach but for LLaMA-2 (Touvron et al., 2023b).
 Jiang et al.
@@
 Wang et al.
-(2023) is concurrent to our work and improves embeddings by adding synthetic training data and trains on Mistral (Jiang et al., 2023a).
-Zero-shot embeddings.
-Most recent sentence embeddings research has focused on improving finetuning.
-Reimers and Gurevych (2019) demonstrates that without finetuning, BERT has low-quality embeddings.
-To our knowledge, Jiang et al.
-(2023b) is the only paper that constructs zero-shot embeddings for autoregressive language models.
+(2023) improves embeddings by adding synthetic training data and trains on Mistral (Jiang et al., 2023a).
+Muennighoff et al.
+(2024) adds an additional objective to maintain generation performace, but requires bidirectional attention.
+Li & Li (2023b) modifies the architecture of a causal language model to be bidirectional.
+Language modeling with bidirectional information.
+Xu et al.
+(2023) finds that repetition can improve the performance of autoregressive language models for reasoning tasks.
+arora2024just finds that repetition can be useful to improve recall for recurrent architectures.
 6 Conclusion
-We have compared classical and echo embeddings in a toy example, on real data in the zero-shot setting, and after finetuning.
-With the toy data, we identified a failure mode of autoregressive classical embeddings, which we have shown can be recovered with echo embeddings.
-Our result motivates the development of higher quality embeddings which are important in retrieval applications.
-In addition, until recently, masked language models largely dominated the MTEB leaderboard, despite often having an order of magnitude fewer parameters, having been trained on substantially less data, and performing worse on other benchmarks of interest to the natural language processing community.
-While our results do not explicitly explain the surprising success of masked language models, they do suggest that next-token language models suffer from an inherent drawback that may have stifled their performance until they became performant enough to compensate for this shortcoming.
-We believe that our embedding strategy achieves the best of both worlds: we gain the capability of next-token language models while recovering from the failure mode that next-token language models do not encode information about future tokens in their contextualized token embeddings.
-8 Limitations
-Despite the success of echo embeddings, the method has limitations.
-First, while echo embeddings achieve superior performance to classical embeddings, they require double the inference cost to pass two copies of the input sequence to the model.
-Though this is double the training cost for a fixed number of training steps, we show in Appendix D that echo embeddings achieve improved performance even when matching compute.
-Second, we do not fully explain why echo embeddings are improved in comparison to classical embeddings after finetuning even though there is no representational limitation.
-We leave it to future work to understand the exact underlying mechanisms for this improvement.
-Acknowledgements
+We have demonstrated that the common assumption that bidirectional architectures are crucial for high quality embedding models is false.
+Our method, echo embeddings, prompts an autoregressive model to act as an autoecoder, which enables the model to encode bidirectional information.
+Our results highlight the importance of bidirectional information—we find that the inability to encode bidirectional information leads to the failure of classical-style embeddings in autoregressive models.
+Further, our method demonstrates that by carefully choosing the task that the language model is prompted to perform at inference time is a promising approach for improving embeddings.
+We hope that our work will inspire future research towards understanding how inference-time changes can impact embeddings.
+Acknowledgments
 This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No.
@@
 In this paper, we use the following models:
-Appendix B Echo Embeddings: Additional Information
+Appendix B Societal Considerations
+In this section we discuss the societal impact of our work and the safeguards we have taken to mitigate negative impacts.
+B.1 Impact
+We intend for and believe that this work will have a positive societal impact.
+In particular, language models are known to suffer from limitations such as hallucinations and biases.
+In the context of retrieval-augmented generation, we believe that higher quality embeddings can lead to more accurate retrievals, which can in turn lead to more accurate generations.
+This can help mitigate the impact of hallucinations and biases in language models.
+We also believe that the improved embeddings can be used in a variety of downstream tasks, such as information retrieval, question answering, and summarization, which can have a positive impact on society.
+Nonetheless, we acknowledge that not all uses of our work will be necessarily positive.
+For example, our work can be used to improve the performance of language models in generating fake news or misinformation.
+Second, we release a training dataset that has been aggregated from a set of other open source datasets that might nonetheless contain harmful content.
+We believe that the positive impacts of our work outweigh the negative impacts, and we are committed to working towards ensuring that our work is used for positive purposes.
+Appendix C Echo Embeddings: Additional Information
 In this section, we aim to describe the additional details that were omitted from Section 2 and 3.
@@
 measures the similarity between the embeddings of $x$ and $y$ for any embedding function $\phi\colon\mathcal{X}\to R^{d}$.
-The cosine similarity is used for our experiments in Sections 3, and as the similarity function for training in 5.
+The cosine similarity is used for our experiments in Sections 3, and as the similarity function for training in 4.4.
 All MTEB datasets use cosine similarity to compute similarity with the exception of the classification datasets, in which similarity is not explicitly measured, and the clustering datasets, which use Euclidean distance,
@@
 Throughout the paper, we use a variety of different prompts to construct embeddings.
-In Section C, we demonstrate that for zero-shot embeddings, the exact wording or template used as a prompting strategy does not have a strong effect on the performance of MTEB tasks, with the exception of for the summarization approach.
+In Section D, we demonstrate that for zero-shot embeddings, the exact wording or template used as a prompting strategy does not have a strong effect on the performance of MTEB tasks, with the exception of for the summarization approach.
 This implies, in general, that classical embeddings and echo embeddings should be robust to the exact choice of prompts.
@@
 For Structure 3, the data is given in Table 10.
-In all cases, the data is generated by GPT4.
-The data from Structure 1 is generated from the following GPT4 prompt, and the other structures are generated from minor variations on this:
-Together, we need to generate sentence triplets. Each triplet will have the following form:
+In all cases, the data is generated by GPT-4.
+The data from Structure 1 is generated from the following GPT-4 prompt, and the other structures are generated from minor variations on this: Together, we need to generate sentence triplets.
+Each triplet will have the following form:
 - Sentence 1 can be anything, be creative here.
@@
 The only difference in wording can be in the second half of the sentence.
-- sentence 3 should be extremely similar to sentence 1 and semantically equivalent, but slightly re-worded
+- Sentence 3 should be extremely similar to sentence 1 and semantically equivalent, but slightly re-worded.
 Here is an example:
-"sentence1": "I like to eat apples and bananas but I really hate almost every other fruit.",
-"sentence2": "I like to eat apples and bananas and I also enjoy also every other fruit",
-"sentence3": "I like two fruits: apples and bananas but I hate nearly all fruits other than these.",
+”sentence1”: ”I like to eat apples and bananas but I really hate almost every other fruit.”,
+”sentence2”: ”I like to eat apples and bananas and I also enjoy also every other fruit”,
+”sentence3”: ”I like two fruits: apples and bananas but I hate nearly all fruits other than these.”,
 The first half of the sentence should be relatively short, less than 10 words, but the second half should be long, at least 10 words. Give more examples, and write them in json format. Be creative!
-Appendix C Additional Zero-shot Results
-In this section, we describe the omitted methodology and results for the zero-shot section.
-Prompt sampling procedure.
-Here we describe the prompt sampling procedure and then provide the prompts that we use for the zero shot:
-  1. 1.
-    Choose an instruction.
-For classical embeddings, we choose from {Write, Say, Complete, Explain}.
-For echo embeddings, we choose from {Repeat, Rewrite, Rephrase, Fill in the blank}.
-For summarization, we choose from {Summarize, Categorize, Understand, Analyze}.
-  2. 2.
-    Choose a wording for the instruction.
-For example, if we chose “Say” as the instruction, then we would choose from {Say a sentence, Say a paragraph, Say something, Say a response, Say a query, Say a prompt}.
-For summarization, we also choose a second part of the wording, as the summarization strategy requires that the summary be in one word: {in one word, with a single word, succinctly with one word, in a unique one-word way, in a single word, in a word}.
-  3. 3.
-    Choose a separator, which include colons, commas, newlines.
-  4. 4.
-    Choose a prefix, which includes markers to indicate the first and appearance of the input.
-  5. 5.
-    Classical prompts have the form: “{instruction} {separator} {prefix} SS”.
-  6. 6.
-[…]
-  7. 7.
-    Summarization prompts have the form: “{instruction0} {separator} {prefix} $S$ {instruction1} {separator}”.
-For classical, we choose the prompts:
-Write a sentence I] S
-Write a prompt!
-(I) S
-Write some text
-PROMPT-S
-SAY A PARAGRAPH | SENTENCE 0] S
-Say a query QUERY: S
-Say a sentence!
-[A] S
-COMPLETE THE PROMPT Text (1) S
-Complete the query SENTENCE 0) S
-Complete the sentence:-S
-Explain a query text 0 S
-Explain a prompt | Sentence 1> S
-EXPLAIN A SENTENCE Prompt (1) S
-For echo, we choose the prompts:
-Repeat The Paragraph.
-query 1) S.
-query 2) S
-Repeat the response.
-1) S.
-AGAIN 2) S
-REPEAT THE SENTENCE :: PROMPT
-S :: RESPONSE
-S
-Rewrite the query | QUERY (A) S | (B) S
-Rewrite the text. SENTENCE A) S. B) S
-Rewrite the response | query A] S | query B] S
-Rephrase the sentence:@S:Again@S
-Rephrase The Sentence!
-Text <> S!
-Answer <> S
-REPHRASE THE QUERY Sentence a) S Answer b) S
-Fill in the blank in the prompt:
-Query a) S:
-Query b) S
-FILL IN THE BLANK IN THE RESPONSE | Sentence A) S | Sentence B) S
-Fill in the blank in the paragraph.
-Text | S.
-Response | S
-For summarization, we use the prompts:
-SUMMARIZE THE QUERY.
-Prompt: SIN A WORD.
-Summarize the sentence!
-PROMPT <1> SSuccinctly With One Word!
-SUMMARIZE THE PARAGRAPH. PROMPT (0) SIN A WORD.
-CATEGORIZE THE PROMPT query
-SWith a single word
-Categorize the query | prompt [1] Sin a word |
-CATEGORIZE THE SENTENCE.
-Prompt <1> SIN A WORD.
-Understand the sentence
-@SIn a single word
-Understand The Prompt:QUERY [0] Sin a single word:}
-UNDERSTAND THE PARAGRAPH:Text I] SSuccinctly with one word:}
-Analyze the sentence.
-Sentence SIn A Unique One-word Way.
-Analyze the response! query a> SIN A UNIQUE ONE-WORD WAY!
-Analyze The Prompt
-Sentence a> SIn a unique one-word way
-Subset of MTEB for zero-shot evaluation.
-We evalaute on the following subset of MTEB: FiQA2018, SCIDOCS, SciFact, NFCorpus, TwitterSemEval2015, TwitterURLCorpus, ImdbClassification, AmazonReviewsClassification, TweetSentimentExtractionClassification, MTOPDomainClassification, TwentyNewsgroupsClustering, BiorxivClusteringS2S, MedrxivClusteringS2S, StackOverflowDupQuestions, AskUbuntuDupQuestions, SciDocsRR, BIOSSES, STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark, and SICK-R.
-Measuring the sensitivity of different embedding strategies to prompting.
-We plot the sensitivity of classical, repetition, and summarization to different choices of prompts for different models in Figures 56, and 7.
-We also extend to plotting on all tested datasets individually in Figures 89, and 10.
-We observe that summarization is highly sensitive to the exact prompt used.
-However, neither classical nor echo were particularly sensitive in any case.
-Consistently, mean token pooling outperformed last token pooling by a large factor.
-Evaluation of zero-shot results with different validation sets.
-We include the zero results of validation using different MTEB datasets.
-For validation, we select one dataset from each category, as follows: classification: ImdbClassification; Pair Classification: TwitterSemEval2015; Clustering: TwentyNewsgroupsClustering; Retrieval: FiQA2018; STS: STSBenchmark, Reranking: StackOverflowDupQuestions.
-We plot these results for different models in Tables 89, and 10.
-We observe similar results across different validation sets, with minor variations in the performance.
-In addition, we the performance of each dataset when the prompts have been validated with FiQA2018 in Tables 1112, and 13.
-Figure 4: We plot the histogram distribution of the difference between the predicted rank and the ground truth rank of sentence pairs in STS datasets.
-When predicted rank is larger than the ground truth rank—when the rank difference is positive—then the embedding has overestimated the similarity of this pair.
-Similarly, negative values imply that the the rank is underestimated.
-We plot the distribution of these ranks for both classical and echo embeddings where we split the data into two groups: one in which sentences are similar in the first part of the sentence (top 10% by first-half similarity), and another in which sentences are similar in the second part of the sentence (top 10% by second-half similarity).
-C.1 Validating the connection between our synthetic data experiments and real data.
-In Section 3, we hypothesized that classical embeddings would overestimate similarity on sentences where the first half of the sentence are similar, and that echo embeddings would recover from this failure mode.
-In order to test this hypothesis, we exact a set of examples from the STS datasets included in the MTEB benchmark in which the first half of the sentence is similar, and measure the degree to which the similarity is overestimated.
-As a control, we also select points which are similar in the second half of the sentence, and measure the degree to which similarity is overestimated.
-By comparing the degree to which sentences which are similar in the first half are overestimated in similarity, and the degree to which sentences which are similar in the second half are overestimated, then we can identify if classical embeddings overestimate similarity in specifically sentences which are similar in the first half.
-Thus, under our hypothesis, we expect that, for classical embeddings, sentences which are similar in the first half are overestimated in similarity more than sentences that are similar in their second half.
-On the other hand, we expect that, for echo embeddings, the degree to which similarity is over- or underestimated is independent of whether the sentences are similar in the first or second half of the sentence.
-Identifying examples based on similarity in the first/second part of the sentence.
-We aim to determine which sentences are most similar in the first half of the sentence or in the second half of the sentence.
-For each sentence pair $x,y$, we split the sentences in half by number of words, yielding $x=[x_{1},x_{2}]$, and $y=[y_{1},y_{2}]$.
-We select sentences which are most similar in the first half by using the off-the-shelf masked-language-model-based embedding model bge-base-en-v1.5 (Xiao et al., 2023b).
-To select sentences that are similar in the first half, we measure the cosine similarity $\operatorname{Sim}(x_{1},y_{1})$ and take the top 10% of sentence pairs $x,y$ which have the highest cosine similarity.
-Similarly, to select sentences which are similar in the second half, we collect the top 10% of examples by $\operatorname{Sim}(x_{2},y_{2})$.
-We collect examples from each of the STS datasets in MTEB.
-Measuring sentence similarity estimation error.
-We must determine the degree to which classical and echo embeddings overestimate similarity.
-The STS datasets contain sentences pairs which are ranked by similarity: the sentences which are most similar have the highest ground-truth ranking, and the least similar sentences have the lowest.
-We will denote the ranking of sentence pair $i$ as $r_{i}$.
-We compute an estimated ranking $\{\hat{r}_{i}\}$ by ranking sentence pairs by the cosine similarity between their embeddings.
-We can compare the error in our estimated ranking by taking the rank difference $\operatorname{Err}_{i}=\hat{r}_{i}-r_{i}$.
-When $\operatorname{Err}_{i}>0$, we say that the $i$th sentence pair is overestimated in similarity, and similarly underestimated when $\operatorname{Err}_{i}<0$.
-Results.
-We plot the the distribution over rank differences for sentences which are similar in the first half and sentences which are similar in the second half for echo and classical embeddings, from all STS datasets (Figure 4).
-We also highlight the means of the distributions.
-In accordance with our hypothesis, we observe that for classical embeddings, sentences which are similar in the first half are generally overestimated in similarity more than sentences which are similar in the second half of the sentence, suggesting that classical embeddings fail particularly on sentences that are similar in early tokens.
-Further, we generally observe no difference between the estimation error distributions for echo embeddings, which demonstrates that echo embeddings recover from this particular failure mode.
-There are some notable counterexamples: BIOSSES does not exhibit this trend, but has few examples and thus the results may arise from noise alone.
-Further, STS22 exhibits identical distributions in estimation error between sentences which are similar in the first half and sentences which are similar in the second half, for both classical and echo embeddings.
-It is unclear why this trend fails to hold for STS22.
-Nonetheless, the trend holds for every other dataset, suggesting that the conceptual failure of classical embeddings that we identified in Section 3 generalizes to real data.
-Qualitative examples.
-In addition, we provide qualitative examples of sentence pairs from STSBenchmark where echo embeddings reduce error most, and where echo embeddings reduce error least, in comparison to classical embeddings.
-More precisely, we plot the top and bottom 7 examples ranked by $|\operatorname{Err}^{\text{classical}}_{i}|-|\operatorname{Err}^{\text{echo}}_{i}|$, where $\operatorname{Err}^{\text{classical}}_{i}$ represents the rank difference of the $i$th example of classical embeddings, and $\operatorname{Err}^{\text{echo}}_{i}$ is similar but for echo embeddings (Table 7).
-Appendix D Additional Finetuning Results
-In this section, we address the omitted details from the finetuning results of the main paper.
-Training Datasets.
-We follow the setup of Wang et al.
-(2023), and use the following datasets: ELI5 (sample ratio 0.1) (Fan et al., 2019), HotpotQA (Yang et al., 2018), FEVER (Thorne et al., 2018), MIRACL (Zhang et al., 2023b), MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) (Bajaj et al., 2018), NQ (Karpukhin et al., 2020), NLI (Gao et al., 2021b), SQuAD (Karpukhin et al., 2020), TriviaQA (Karpukhin et al., 2020), Quora Duplicate Questions (sample ratio 0.1) (DataCanary et al., 2017), Mr- TyDi (Zhang et al., 2021), DuReader (Qiu et al., 2022), and T2Ranking (sample ratio 0.5) (Xie et al., 2023).
-We use approximately 1.5M training examples.
-GPUs.
-Training a model takes approximately two days on 4 A100 GPUs.
-Instructions for finetuning datasets.
-We also follow the setup of Wang et al. (2023), and use the instructions in Table 3. For evaluation, we use the instructions found in Table 14.
-Models on MTEB leaderboard.
-We compare our implementation of classical and echo embeddings to state-of-the-art approaches on MTEB.
-Namely, we display results for UAE-Large-V1 (Li and Li, 2023), multilingual-e5-large (Wang et al., 2024), bge-large-en-v1.5 (Xiao et al., 2023b), udever-bloom-7b (Zhang et al., 2023a), sgpt-5.8b (Muennighoff, 2022), e5-mistral-7b (concurrent work) (Wang et al., 2023).
-Additional ablations.
-We plot additional ablations, including ablating the role of instructions during training and evaluation, as well as providing an evaluation at step 280 (out of 720 total steps), which is approximately $1/3$ of the duration of training (Table 15).
-We note that echo embeddings still outperform classical embeddings in this setting.
-Performance over training time.
-We plot the performance over the duration of training for a subset of MTEB tasks in Figure 11. Surprisingly, task performance decreases over training for many tasks.
-Computational benefits of echo embeddings.
-From Table 15, we observe that even after approximately $1/3$ of the total training duration (less than $1/2$), echo embeddings achieve performance higher than classical embeddings achieve after an entire epoch (Table 2).
-Echo embeddings requires twice the computational cost of classical embeddings.
-However, this result suggests that despite this additional cost per embedding, training with echo embeddings can save on training costs by requiring less than half an epoch of training to outperform classical embeddings.
-Further, since each data point is only seen once, it implies that echo embeddings are much more data efficient than classical embeddings, which may be helpful when data is costly or difficult to acquire.
-All results.
-We plot the results for every MTEB dataset for echo embeddings, for classical embeddings, and for bidirectional embeddings in Table 16.
-NLI Given a premise, retrieve a hypothesis that is entailed by the premise
-NLI Retrieve semantically similar text
-DuReader Given a Chinese search query, retrieve web passages that answer the question
-ELI5 Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
-FEVER Given a claim, retrieve documents that support or refute the claim
-HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
-MIRACL Given a question, retrieve Wikipedia passages that answer the question
-MrTyDi Given a question, retrieve Wikipedia passages that answer the question
-MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query
-MSMARCO Document Given a web search query, retrieve relevant documents that answer the query
-NQ Given a question, retrieve Wikipedia passages that answer the question
-QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question
-QuoraDuplicates Find questions that have the same meaning as the input question
-Squad Retrieve Wikipedia passages that answer the question
-T2Ranking Given a Chinese search query, retrieve web passages that answer the question
-TriviaQA Retrieve Wikipedia passages that answer the question
-Table 3: Instructions for finetuning datasets.
-Training objective.
-For the training objective, we use the SimCSE loss (Gao et al., 2021b). It is defined,
-$\ell_{i}=-\log\frac{\exp\left(\operatorname{Sim}\left(h_{i},h_{i}^{+}\right)/\tau\right)}{\sum_{j=1}^{N}\exp(\operatorname{Sim}\left(h_{i},h_{j}^{-}\right)/\tau)}.$ (3)
-In this loss function, $h_{i}$ represents a query (or a reference sentence when the data is symmetric), $h_{i}^{+}$ represents a positive example associated with $h_{i}$, and $\{h_{j}^{-}\}_{j=1}^{N}$ represents the set of negatives associated with the example, including mined hard negatives.
 $q$ $^{-} s+s^{+}$
@@
 3.6
-What kind of insulation is it? What kind of floors are above? 0.0 "Tony’s not feeling well," Spurs coach Gregg Popovich said. We’re thrilled to be up 3-2,” Coach Gregg Popovich said Wednesday. 1.6
+What kind of insulation is it? What kind of floors are above? 0.0 ”Tony’s not feeling well,” Spurs coach Gregg Popovich said. We’re thrilled to be up 3-2,” Coach Gregg Popovich said Wednesday. 1.6
 It depends entirely on your company and your contract.
@@
 Table 11: Example sentences from STSBenchmark in which zero-shot echo embeddings with Mistral 7B most improve (left) and least improve (right).
+Appendix D Additional Zero-shot Results
+Recent literature suggests that the performance of language models on zero-shot tasks can be highly variable depending on the exact wording and template of the prompts (Sclar et al., 2023).
+Thus, in order to test the degree to which the exact prompt matters for each of the embedding strategies we consider, we perform prompt randomization where we sample prompts by randomizing the exact wording, punctuation, and capitalization of the prompt.
+Prompt sampling procedure.
+Here we describe the prompt sampling procedure and then provide the prompts that we use for the zero shot:
+  1. 1.
+    Choose an instruction.
+For classical embeddings, we choose from {Write, Say, Complete, Explain}.
+For echo embeddings, we choose from {Repeat, Rewrite, Rephrase, Fill in the blank}.
+For PromptEOL, we choose from {Summarize, Categorize, Understand, Analyze}.
+  2. 2.
+    Choose a wording for the instruction.
+For example, if we chose “Say” as the instruction, then we would choose from {Say a sentence, Say a paragraph, Say something, Say a response, Say a query, Say a prompt}.
+For PromptEOL, we also choose a second part of the wording, as the PromptEOL strategy requires that the summary be in one word: {in one word, with a single word, succinctly with one word, in a unique one-word way, in a single word, in a word}.
+  3. 3.
+    Choose a separator, which include colons, commas, newlines.
+  4. 4.
+    Choose a prefix, which includes markers to indicate the first and appearance of the input.
+  5. 5.
+    Classical prompts have the form: “{instruction} {separator} {prefix} SS”.
+  6. 6.
+[…]
+  7. 7.
+    PromptEOL prompts have the form: “{instruction0} {separator} {prefix} $S$ {instruction1} {separator}”.
+Classical embeddings.
+For classical, we choose the prompts:
+Write a sentence I] $S$
+Write a prompt!
+(I) $S$
+Write some text
+PROMPT-$S$
+SAY A PARAGRAPH — SENTENCE 0] $S$
+Say a query QUERY: $S$
+Say a sentence!
+[A] $S$
+COMPLETE THE PROMPT Text (1) $S$
+COMPLETE THE QUERY text 0) $S$
+Complete the query SENTENCE 0) $S$
+Complete the sentence:-$S$
+Explain a query text 0 $S$
+Explain a prompt — Sentence 1¿ $S$
+EXPLAIN A SENTENCE Prompt (1) $S$
+Echo embeddings.
+For echo, we choose the prompts:
+Repeat The Paragraph.
+query 1) $S$.
+query 2) $S^{\prime}$
+Repeat the response.
+1) $S$.
+AGAIN 2) $S^{\prime}$
+REPEAT THE SENTENCE :: PROMPT
+$S$ :: RESPONSE
+S’
+Rewrite the query — QUERY (A) $S$ — (B) $S^{\prime}$
+Rewrite the text. SENTENCE A) $S$. B) $S^{\prime}$
+Rewrite the response — query A] $S$ — query B] $S^{\prime}$
+Rephrase the sentence:@$S$:Again@S’
+Rephrase The Sentence! Text ¡¿ SS! Answer ¡¿ $S^{\prime}$
+REPHRASE THE QUERY Sentence a) $S$ Answer b) $S^{\prime}$
+Fill in the blank in the prompt: Query a) $S$: Query b) $S^{\prime}$
+FILL IN THE BLANK IN THE RESPONSE — Sentence A) $S$ — Sentence B) $S^{\prime}$
+Fill in the blank in the paragraph.
+Text — $S$.
+Response — $S^{\prime}$
+PromptEOL embeddings.
+For PromptEOL, we use the prompts:
+SUMMARIZE THE QUERY.
+Prompt: $S$’IN A WORD.
+Summarize the sentence!
+PROMPT ¡1¿ $S$’Succinctly With One Word!
+SUMMARIZE THE PARAGRAPH. PROMPT (0) $S$’IN A WORD.
+CATEGORIZE THE PROMPT query
+$S$ With a single word
+Categorize the query — prompt [1] $S$ in a word —
+CATEGORIZE THE SENTENCE.
+Prompt ¡1¿ $S$ IN A WORD.
+Understand the sentence
+@$S$ In a single word
+Understand The Prompt:QUERY [0] $S$ in a single word:
+UNDERSTAND THE PARAGRAPH:Text I] $S$ Succinctly with one word:
+Analyze the sentence.
+Sentence $S$ In A Unique One-word Way.
+Analyze the response! query a¿ $S$ IN A UNIQUE ONE-WORD WAY!
+Analyze The Prompt
+Sentence a¿ $S$ In a unique one-word way
+Subset of MTEB for zero-shot sensitivity evaluation.
+Given that we evaluate on hundreds of prompts, we choose a subset of MTEB to reflect each of the MTEB categories but with a substantially reduced computational requirement for evaluation.
+We evalaute on the following subset of MTEB: FiQA2018, SCIDOCS, SciFact, NFCorpus, TwitterSemEval2015, TwitterURLCorpus, ImdbClassification, AmazonReviewsClassification, TweetSentimentExtractionClassification, MTOPDomainClassification, TwentyNewsgroupsClustering, BiorxivClusteringS2S, MedrxivClusteringS2S, StackOverflowDupQuestions, AskUbuntuDupQuestions, SciDocsRR, BIOSSES, STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark, and SICK-R.
+Note that this excludes the largest retrieval datasets in MTEB, on which we find PromptEOL is most likely to fail.
+Measuring the sensitivity of different embedding strategies to prompting.
+We plot the sensitivity of classical, repetition, and PromptEOL to different choices of prompts for different models in Figures 78, and 9.
+We also extend to plotting on all tested datasets individually in Figures 1011, and 12.
+We observe that PromptEOL is highly sensitive to the exact prompt used.
+However, neither classical nor echo were particularly sensitive.
+Consistently, mean token pooling outperformed last token pooling by a large factor.
+Figure 6: We plot the histogram distribution of the difference between the predicted rank and the ground truth rank of sentence pairs in STS datasets.
+When predicted rank is larger than the ground truth rank—when the rank difference is positive—then the embedding has overestimated the similarity of this pair.
+Similarly, negative values imply that the the rank is underestimated.
+We plot the distribution of these ranks for both classical and echo embeddings where we split the data into two groups: one in which sentences are similar in the first part of the sentence (top 10% by first-half similarity), and another in which sentences are similar in the second part of the sentence (top 10% by second-half similarity).
+Appendix E Toy Verification
+In Section 3, we demonstrate that classical embeddings overestimate the similarity between examples which are superficially similar based on tokens that appear early in the sequence.
+In this section, we investigate whether this failure mode generalizes to real data.
+We provide qualitative and quantitative evidence that this failure mode does, in fact, generalize to real data.
+E.1 Qualitative comparison of classical and echo embeddings.
+To build intuition that this applies to realistic data, we present the sentence pair from STSBenchmark, a sentence similarity task from MTEB, in which echo embeddings reduce error the most:
+[…]
+which has a ground-truth score of 0 (out of $5$) similarity. The sentence pair for which echo embeddings reduces error the least is:
+[…]
+which has a ground-truth similarity of $2$ (out of $5$). We provide more examples in the Appendix Table 11.
+For these examples, the sentence pair $(x_{1},x_{2})$ on which echo embeddings improve error the most has exactly the property we identify as a failure mode for classical embeddings: the sentence is superficially similar for the first few tokens.
+On the other hand $(y_{1},y_{2})$ does not have this property.
+E.2 Quantitative evaluation of the failure mode.
+In order to test this hypothesis quantitatively, we design an experiment to identify real examples that are superficially similar in their early tokens, and examples that do not have this property.
+We compare the behavior of classical embeddings on these datapoints.
+To conduct this experiment, we extract a set of examples from the STS datasets included in the MTEB benchmark in which the first half of the sentence is similar, and measure the degree to which the similarity is overestimated.
+We also select points which are similar in the second half of the sentence, and measure the degree to which similarity is overestimated.
+By comparing the degree to which sentences which are similar in the first half are overestimated in similarity, and the degree to which sentences which are similar in the second half are overestimated, then we can identify if classical embeddings overestimate similarity in specifically sentences which are similar in the first half.
+Thus, under our hypothesis, we expect that, for classical embeddings, sentences which are similar in the first half are overestimated in similarity more than sentences that are similar in their second half.
+On the other hand, we expect that, for echo embeddings, the degree to which similarity is over- or underestimated is independent of whether the sentences are similar in the first or second half of the sentence.
+Identifying examples based on similarity in the first/second part of the sentence.
+We aim to determine which sentences are most similar in the first half of the sentence or in the second half of the sentence.
+For each sentence pair $x,y$, we split the sentences in half by number of words, yielding $x=[x_{1},x_{2}]$, and $y=[y_{1},y_{2}]$.
+We select sentences which are most similar in the first half by using the off-the-shelf masked-language-model-based embedding model bge-base-en-v1.5 (Xiao et al., 2023b).
+To select sentences that are similar in the first half, we measure the cosine similarity $\operatorname{Sim}(x_{1},y_{1})$ and take the top 10% of sentence pairs $x,y$ which have the highest cosine similarity.
+Similarly, to select sentences which are similar in the second half, we collect the top 10% of examples by $\operatorname{Sim}(x_{2},y_{2})$.
+We collect examples from each of the STS datasets in MTEB.
+Measuring sentence similarity estimation error.
+We must determine the degree to which classical and echo embeddings overestimate similarity.
+The STS datasets contain sentences pairs which are ranked by similarity: the sentences which are most similar have the highest ground-truth ranking, and the least similar sentences have the lowest.
+We will denote the ranking of sentence pair $i$ as $r_{i}$.
+We compute an estimated ranking $\{\hat{r}_{i}\}$ by ranking sentence pairs by the cosine similarity between their embeddings.
+We can compare the error in our estimated ranking by taking the rank difference $\operatorname{Err}_{i}=\hat{r}_{i}-r_{i}$.
+When $\operatorname{Err}_{i}>0$, we say that the $i$th sentence pair is overestimated in similarity, and similarly underestimated when $\operatorname{Err}_{i}<0$.
+Results.
+We plot the the distribution over rank differences for sentences which are similar in the first half and sentences which are similar in the second half for echo and classical embeddings, from all STS datasets (Figure 6).
+We also highlight the means of the distributions.
+In accordance with our hypothesis, we observe that for classical embeddings, sentences which are similar in the first half are generally overestimated in similarity more than sentences which are similar in the second half of the sentence, suggesting that classical embeddings fail particularly on sentences that are similar in early tokens.
+Further, we generally observe no difference between the estimation error distributions for echo embeddings, which demonstrates that echo embeddings recover from this particular failure mode.
+There are some notable counterexamples: BIOSSES does not exhibit this trend, but has few examples and thus the results may arise from noise alone.
+Further, STS22 exhibits identical distributions in estimation error between sentences which are similar in the first half and sentences which are similar in the second half, for both classical and echo embeddings.
+It is unclear why this trend fails to hold for STS22.
+Nonetheless, the trend holds for every other dataset, suggesting that the conceptual failure of classical embeddings that we identified in Section 3 generalizes to real data.
+Qualitative examples.
+In addition, we provide qualitative examples of sentence pairs from STSBenchmark where echo embeddings reduce error most, and where echo embeddings reduce error least, in comparison to classical embeddings.
+More precisely, we plot the top and bottom 7 examples ranked by $|\operatorname{Err}^{\text{classical}}_{i}|-|\operatorname{Err}^{\text{echo}}_{i}|$, where $\operatorname{Err}^{\text{classical}}_{i}$ represents the rank difference of the $i$th example of classical embeddings, and $\operatorname{Err}^{\text{echo}}_{i}$ is similar but for echo embeddings (Table 11).
+E.3 Full zero-shot
+Here we present omitted results from the main paper. First, the role of scale and base model is shown in Table 12
+Table 12: Role of scale and base model in the zero-shot setting: We study the role of scale and base model using a different 7B parameter model: LLaMA-7B and a 1.3B parameter model: S-LLaMA-1.3B.
+Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
+# of datasets 12 11 3 4 15 10 1 56
+Echo 1.3B 64.85 30.62 47.30 42.35 13.23 61.68 23.80 40.45
+Classical 1.3B 59.41 27.50 41.01 38.09 8.53 47.91 23.36 34.31
+PromptEOL 1.3B 68.86 24.52 36.52 43.81 8.34 56.64 27.36 37.50
+Echo 7B 70.75 30.68 65.15 45.72 18.03 70.19 29.81 45.84
+Classical 7B 63.54 31.99 53.78 42.82 15.07 55.69 26.21 40.29
+PromptEOL 7B 74.02 26.93 58.04 47.48 10.11 66.94 29.71 42.84
 Figure 7: Variance over different prompting strategies for zero-shot Mistral-7B.
@@
 Figure 12: Variance over different prompting strategies for all evaluated datasets for zero-shot LLaMa-2-13B.
-Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
-Classical
-Classification 59.20 73.80 24.16 20.57 58.59 54.54 46.79
-Pair Classification 58.73 71.40 24.32 20.39 59.00 54.42 46.64
-Clustering 58.23 72.62 23.90 18.64 56.68 54.82 45.37
-Retrieval 58.21 73.87 23.85 20.35 56.97 54.44 45.88
-STS 58.31 44.03 13.07 2.63 38.95 46.77 34.63
-Reranking 58.13 71.77 24.20 20.23 58.59 54.89 46.43
-Echo
-Classification 64.50 74.65 25.93 22.52 73.81 59.41 55.57
-Pair Classification 64.15 75.93 22.25 18.35 72.75 58.47 54.15
-Clustering 61.54 71.04 26.32 15.88 68.18 60.27 51.81
-Retrieval 64.06 75.26 27.02 23.61 72.40 60.00 55.07
-STS 64.50 74.65 25.93 22.52 73.81 59.41 55.57
-Reranking 64.15 75.93 22.25 18.35 72.75 58.47 54.15
-Summarization
-Classification 66.62 78.95 21.79 14.68 72.13 64.24 55.22
-Pair Classification 66.62 78.95 21.79 14.68 72.13 64.24 55.22
-Clustering 66.66 79.59 28.08 11.88 67.30 65.19 53.43
-Retrieval 66.01 81.82 26.48 19.13 70.13 66.24 54.96
-STS 66.01 81.82 26.48 19.13 70.13 66.24 54.96
-Reranking 63.19 75.22 26.09 20.52 65.98 59.05 51.55
-Table 8: Scores for additional zero-shot validation datasets on Mistral-7B.
-Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
-Classical
-Classification 57.59 68.65 23.72 18.06 57.19 54.59 45.14
-Pair Classification 57.56 70.18 23.51 18.54 58.24 54.40 45.79
-Clustering 57.14 69.91 23.35 16.98 57.66 55.38 45.25
-Retrieval 56.61 68.46 23.22 18.63 56.49 53.26 44.65
-STS 57.56 70.18 23.51 18.54 58.24 54.40 45.79
-Reranking 56.65 66.54 22.46 10.48 55.97 54.44 42.98
-Echo
-Classification 62.24 67.96 23.60 14.33 65.79 55.44 49.85
-Pair Classification 63.42 72.52 21.11 17.35 68.16 54.98 51.47
-Clustering 60.12 66.74 23.45 11.60 64.45 56.31 48.75
-Retrieval 61.64 66.29 25.11 16.12 66.18 56.35 50.26
-STS 63.15 68.74 23.65 16.38 69.37 57.75 51.96
-Reranking 62.30 74.23 24.69 18.17 65.07 56.76 50.51
-Summarization
-Classification 63.96 77.93 21.89 15.93 67.07 63.39 52.34
-Pair Classification 63.96 77.93 21.89 15.93 67.07 63.39 52.34
-Clustering 61.60 69.47 24.44 5.28 57.53 57.62 45.85
-Retrieval 64.90 78.74 26.63 15.59 70.15 65.43 54.02
-STS 64.90 78.74 26.63 15.59 70.15 65.43 54.02
-Reranking 60.54 69.73 26.40 15.82 61.60 58.80 47.83
-Table 9: Scores for additional zero-shot validation datasets on LLaMa-2-7B.
-Validation Dataset Classification Pair Classification Clustering Retrieval STS Reranking Average
-Classical
-Classification 58.24 71.65 23.91 21.79 58.74 56.37 46.66
-Pair Classification 58.10 73.30 23.01 16.97 57.83 56.17 45.52
-Clustering 58.61 67.47 23.30 15.51 57.93 56.86 45.05
-Retrieval 58.50 65.06 24.22 18.92 57.47 56.38 45.15
-STS 58.24 71.65 23.91 21.79 58.74 56.37 46.66
-Reranking 58.61 67.47 23.30 15.51 57.93 56.86 45.05
-Echo
-Classification 64.15 74.22 25.02 27.58 70.81 61.43 55.02
-Pair Classification 64.57 77.63 22.56 24.08 69.85 59.89 53.55
-Clustering 63.26 73.50 25.10 27.48 69.04 61.81 54.32
-Retrieval 64.65 74.57 25.72 26.58 72.20 62.68 55.60
-STS 63.16 75.98 24.08 27.56 71.00 61.84 54.85
-Reranking 62.90 70.58 25.53 22.11 68.82 62.38 53.02
-Summarization
-Classification 66.02 79.06 26.47 22.20 67.91 64.90 54.52
-Pair Classification 66.02 79.06 26.47 22.20 67.91 64.90 54.52
-Clustering 63.84 71.98 21.99 7.48 56.96 59.41 46.50
-Retrieval 66.02 79.06 26.47 22.20 67.91 64.90 54.52
-STS 66.02 79.06 26.47 22.20 67.91 64.90 54.52
-Reranking 61.19 69.63 26.38 19.62 60.76 62.79 48.36
-Table 10: Scores for additional zero-shot validation datasets on LLaMa-2-13B.
-Dataset Classical Echo Summarization
-FiQA2018 (retrieval) 7.89 12.74 12.43
-SCIDOCS (retrieval) 3.60 4.88 9.97
-SciFact (retrieval) 45.39 49.36 29.90
-NFCorpus (retrieval) 12.07 16.57 17.51
-TwitterSemEval20. (pair_classification) 47.81 62.49 59.79
-TwitterURLCorpus (pair_classification) 73.87 75.26 81.82
-ImdbClassificati. (classification) 72.50 72.02 82.78
-AmazonReviewsCla. (classification) 37.09 40.72 45.58
-TweetSentimentEx. (classification) 53.70 58.76 61.74
-MTOPDomainClassi. (classification) 83.85 92.71 90.72
-TwentyNewsgroups. (clustering) 20.84 29.48 30.11
-BiorxivClusterin. (clustering) 23.47 27.61 27.21
-MedrxivClusterin. (clustering) 24.23 26.42 25.75
-StackOverflowDup. (reranking) 35.85 42.71 40.32
-AskUbuntuDupQues. (reranking) 49.49 54.09 57.17
-SciDocsRR (reranking) 59.38 65.91 75.30
-BIOSSES (sts) 59.05 78.19 66.06
-STS12 (sts) 42.01 58.43 64.62
-STS13 (sts) 59.66 78.53 78.45
-STS14 (sts) 50.69 68.42 71.00
-STS15 (sts) 61.81 78.82 78.29
-STS16 (sts) 57.03 77.52 77.40
-STS17 (sts) 68.08 82.14 78.80
-STS22 (sts) 61.23 57.60 47.07
-STSBenchmark (sts) 47.55 73.85 77.39
-SICK-R (sts) 53.19 71.95 69.48
-Average 45.88 55.07 54.96
-Table 11: Evaluation of all MTEB datasets for zero-shot for Mistral-7B.
-Dataset Classical Echo Summarization
-FiQA2018 (retrieval) 6.48 12.38 9.00
-SCIDOCS (retrieval) 3.72 4.38 8.33
-SciFact (retrieval) 42.18 30.61 23.01
-NFCorpus (retrieval) 10.01 13.38 15.43
-TwitterSemEval20. (pair_classification) 44.11 54.66 54.27
-TwitterURLCorpus (pair_classification) 68.46 66.29 78.74
-ImdbClassificati. (classification) 71.65 73.11 85.83
-AmazonReviewsCla. (classification) 36.16 40.68 44.77
-TweetSentimentEx. (classification) 52.04 54.85 59.96
-MTOPDomainClassi. (classification) 81.63 89.38 89.97
-TwentyNewsgroups. (clustering) 15.88 23.42 32.28
-BiorxivClusterin. (clustering) 23.13 25.92 27.79
-MedrxivClusterin. (clustering) 23.31 24.30 25.48
-StackOverflowDup. (reranking) 35.57 40.82 35.63
-AskUbuntuDupQues. (reranking) 48.51 51.42 56.09
-SciDocsRR (reranking) 58.01 61.29 74.76
-BIOSSES (sts) 65.31 71.96 68.04
-STS12 (sts) 41.84 52.40 60.20
-STS13 (sts) 58.43 72.40 76.31
-STS14 (sts) 49.21 61.24 68.73
-STS15 (sts) 60.03 72.67 75.59
-STS16 (sts) 56.40 73.51 76.71
-STS17 (sts) 62.31 71.87 79.38
-STS22 (sts) 59.48 55.21 55.69
-STSBenchmark (sts) 49.45 65.73 76.42
-SICK-R (sts) 55.35 64.39 70.69
-Average 44.65 50.26 54.02
-Table 12: Evaluation of all MTEB datasets for zero-shot for LLaMa-2-7B.
-Dataset Classical Echo Summarization
-FiQA2018 (retrieval) 8.31 18.07 9.43
-SCIDOCS (retrieval) 4.87 7.56 10.38
-SciFact (retrieval) 41.64 50.55 40.19
-NFCorpus (retrieval) 10.26 21.63 16.02
-TwitterSemEval20. (pair_classification) 42.43 62.85 59.55
-TwitterURLCorpus (pair_classification) 65.06 74.57 79.06
-ImdbClassificati. (classification) 71.82 75.44 91.86
-AmazonReviewsCla. (classification) 37.88 43.25 50.60
-TweetSentimentEx. (classification) 52.95 58.18 59.93
-MTOPDomainClassi. (classification) 84.67 92.52 87.51
-TwentyNewsgroups. (clustering) 17.21 25.98 32.08
-BiorxivClusterin. (clustering) 24.95 26.75 28.30
-MedrxivClusterin. (clustering) 23.49 24.70 24.64
-StackOverflowDup. (reranking) 37.24 44.86 38.44
-AskUbuntuDupQues. (reranking) 50.74 55.21 54.15
-SciDocsRR (reranking) 62.03 70.15 75.65
-BIOSSES (sts) 63.26 77.60 69.33
-STS12 (sts) 51.80 59.36 51.17
-STS13 (sts) 61.59 79.01 76.08
-STS14 (sts) 49.69 69.75 66.62
-STS15 (sts) 58.48 79.86 73.75
-STS16 (sts) 53.18 76.75 77.40
-STS17 (sts) 65.10 80.41 75.88
-STS22 (sts) 59.00 56.84 49.23
-STSBenchmark (sts) 44.80 71.31 75.17
-SICK-R (sts) 55.13 70.27 71.70
-Average 45.15 55.60 54.52
-Table 13: Evaluation of all MTEB datasets for zero-shot for LLaMa-2-13B.
+Appendix F Additional Finetuning Results
+Table 13: MTEB Leaderboard: We compare to other recent models on MTEB. For a more complete list, please visit: https://huggingface.co/spaces/mteb/leaderboard.
+Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
+# of datasets 12 11 3 4 15 10 1 56
+Main fine-tuning results trained with publicly available data:
+Echo (ours) 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
+Echo + compute matched 77.39 46.27 87.49 58.08 55.07 83.18 30.73 64.66
+Classical 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
+Classical + bidirectional attn. 76.70 45.94 88.15 57.23 54.96 82.42 29.32 64.23
+UAE-Large-V1 (MLM) 75.58 46.73 87.25 59.88 54.66 84.54 32.03 64.64
+multilingual-e5-large (MLM) 77.56 47.10 86.19 58.58 52.47 84.78 30.39 64.41
+bge-large-en-v1.5 (MLM) 75.97 46.08 87.12 60.03 54.29 83.11 31.61 64.23
+udever-bloom-7b (autoregr.) 72.13 40.81 85.40 55.91 49.34 83.01 30.97 60.63
+sgpt-5.8b (autoregr.) 68.13 40.34 82.00 56.56 50.25 78.10 31.46 58.93
+Prior work trained with closed-source data (see Section 5.2):
+e5-mistral-7b (autoregr.) 78.47 50.26 88.34 60.21 56.89 84.63 31.40 66.63
+GritLM-7B (autoregr.) 79.46 50.61 87.16 60.49 57.41 83.35 30.37 66.76
+In this section, we address the omitted details from the finetuning results of the main paper.
+Training Datasets.
+We follow the setup of Wang et al.
+(2023), and use the following datasets: ELI5 (sample ratio 0.1) (Fan et al., 2019), HotpotQA (Yang et al., 2018), FEVER (Thorne et al., 2018), MIRACL (Zhang et al., 2023b), MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) (Bajaj et al., 2018), NQ (Karpukhin et al., 2020), NLI (Gao et al., 2021b), SQuAD (Karpukhin et al., 2020), TriviaQA (Karpukhin et al., 2020), Quora Duplicate Questions (sample ratio 0.1) (DataCanary et al., 2017), Mr- TyDi (Zhang et al., 2021), DuReader (Qiu et al., 2022), and T2Ranking (sample ratio 0.5) (Xie et al., 2023).
+We use approximately 1.5M training examples.
+GPUs.
+Training a model takes approximately two days on four A100 GPUs. Evaluating a model on the MTEB benchmark can be completed in parallel in approximately two days with eight A100s.
+Instructions for finetuning datasets.
+We also follow the setup of Wang et al. (2023), and use the instructions in Table 15. For evaluation, we use the instructions found in Table 16.
+Models on MTEB leaderboard.
+We compare our implementation of classical and echo embeddings to state-of-the-art approaches on MTEB.
+Namely, we display results for UAE-Large-V1 (Li & Li, 2023a), multilingual-e5-large (Wang et al., 2024), bge-large-en-v1.5 (Xiao et al., 2023b), udever-bloom-7b (Zhang et al., 2023a), sgpt-5.8b (Muennighoff, 2022), e5-mistral-7b (concurrent work) (Wang et al., 2023).
+Pooling.
+We investigate the role of pooling for fine-tuning (Table 14). We find that last-token, with a trainable end-of-sentence token embedding, is optimal.
+Table 14: Role of pooling in the fine-tuning setting: Average MTEB Score (56 datasets) for different pooling strategies.
+We evaluate mean and last-token pooling after fine-tuning with Mistral-7B-Instruct-v0.1.
+Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Average
+# of datasets 12 11 3 4 15 10 1 56
+Echo + last token pooling 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
+Echo + mean pooling 77.00 44.94 87.73 58.30 55.11 82.52 29.46 64.22
+Classical + last token pooling 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
+Classical + mean pooling 76.26 42.68 86.31 57.58 53.75 81.53 30.19 62.96
+Additional ablations.
+We plot additional ablations, including ablating the role of instructions during training and evaluation, as well as providing an evaluation at step 280 (out of 720 total steps), which is approximately $40\%$ of the duration of training (Table 17).
+We note that echo embeddings still outperform classical embeddings in this setting.
+Comparisons to the MTEB leaderboard.
+We include some recent models on the MTEB leaderboard in Table 13.
+Performance over training time.
+We plot the performance over the duration of training for a subset of MTEB tasks in Figure 13. Surprisingly, task performance decreases over training for many tasks.
+All results.
+We plot the results for every MTEB dataset for echo embeddings, for classical embeddings, and for bidirectional embeddings in Table 18.
+NLI Given a premise, retrieve a hypothesis that is entailed by the premise
+NLI Retrieve semantically similar text
+DuReader Given a Chinese search query, retrieve web passages that answer the question
+ELI5 Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
+FEVER Given a claim, retrieve documents that support or refute the claim
+HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
+MIRACL Given a question, retrieve Wikipedia passages that answer the question
+MrTyDi Given a question, retrieve Wikipedia passages that answer the question
+MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query
+MSMARCO Document Given a web search query, retrieve relevant documents that answer the query
+NQ Given a question, retrieve Wikipedia passages that answer the question
+QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question
+QuoraDuplicates Find questions that have the same meaning as the input question
+Squad Retrieve Wikipedia passages that answer the question
+T2Ranking Given a Chinese search query, retrieve web passages that answer the question
+TriviaQA Retrieve Wikipedia passages that answer the question
+Table 15: Instructions for finetuning datasets.
+Training objective.
+For the training objective, we use the SimCSE loss (Gao et al., 2021b). It is defined,
+$\ell_{i}=-\log\frac{\exp\left(\operatorname{Sim}\left(h_{i},h_{i}^{+}\right)/\tau\right)}{\sum_{j=1}^{N}\exp(\operatorname{Sim}\left(h_{i},h_{j}^{-}\right)/\tau)}.$ (3)
+In this loss function, $h_{i}$ represents a query (or a reference sentence when the data is symmetric), $h_{i}^{+}$ represents a positive example associated with $h_{i}$, and $\{h_{j}^{-}\}_{j=1}^{N}$ represents the set of negatives associated with the example, including mined hard negatives.
 AmazonCounterfactualCls. Classify a given Amazon customer review text as either counterfactual or not counterfactual
@@
 Figure 13: Performance of the evaluated MTEB datasets for finetuning over the number of finetuning steps.
-Model Average Clas. Clus. Pair Clas. Rera. Retr. STS Summ.
-Classical (w/ instruct., mean) 62.96 76.26 42.68 86.31 57.58 53.75 81.53 30.19
-Classical (w/ instruct., last) 63.98 76.57 45.78 86.37 56.71 54.87 82.03 31.02
-Echo (w/ instruct., mean) 64.22 77.00 44.94 87.73 58.30 55.11 82.52 29.46
-Echo (w/ instruct., last) 64.68 77.43 46.32 87.34 58.14 55.52 82.56 30.73
-Classical (w/out instruct., mean) 62.19 75.23 41.79 85.24 56.31 53.24 80.97 30.64
-Classical (w/out instruct., last) 62.37 75.01 42.70 85.69 56.64 53.29 80.92 30.91
-Echo (w/out instruct., mean) 63.28 75.26 42.93 86.95 57.05 55.65 81.40 30.62
-Echo (w/out instruct., last) 62.80 75.30 42.94 86.31 57.31 54.18 80.92 31.00
-Classical (w/ instruct., mean, step 280) 63.19 76.18 42.99 85.44 57.63 53.96 82.53 29.94
-Classical (w/ instruct., last, step 280) 63.87 76.54 46.22 86.70 57.79 53.73 82.22 30.13
-Echo (w/ instruct., mean, step 280) 64.04 76.84 45.76 87.72 59.33 53.55 82.64 30.33
-Echo (w/ instruct., last, step 280) 64.50 76.41 46.70 87.17 59.10 54.84 82.98 31.09
 Table 17: Additional ablations for finetuning.
-Dataset Repetition (last) Repetition (mean) Classical (last) Classical (mean) Bidirectional (last)
+All ablations were performed using Mistral-7B as a backbone.
+We include ablations to compare the effect of using instructions, the effect of using the mean or last token, and the effect of using the last token at a different step.
+Categories $\longrightarrow$ Clas. Clus. P. Cls. Rera. Retr. STS Su. Avg
+# of datasets 12 11 3 4 15 10 1 56
+Echo (w/ instruct., mean) 77.00 44.94 87.73 58.30 55.11 82.52 29.46 64.22
+Echo (w/ instruct., last) 77.43 46.32 87.34 58.14 55.52 82.56 30.73 64.68
+Classical (w/ instruct., mean) 76.26 42.68 86.31 57.58 53.75 81.53 30.19 62.96
+Classical (w/ instruct., last) 76.57 45.78 86.37 56.71 54.87 82.03 31.02 63.98
+Echo (w/out instruct., mean) 75.26 42.93 86.95 57.05 55.65 81.40 30.62 63.28
+Echo (w/out instruct., last) 75.30 42.94 86.31 57.31 54.18 80.92 31.00 62.80
+Classical (w/out instruct., mean) 75.23 41.79 85.24 56.31 53.24 80.97 30.64 62.19
+Classical (w/out instruct., last) 75.01 42.70 85.69 56.64 53.29 80.92 30.91 62.37
+Echo (w/ instruct., mean, step 280) 76.84 45.76 87.72 59.33 53.55 82.64 30.33 64.04
+Echo (w/ instruct., last, step 280) 76.41 46.70 87.17 59.10 54.84 82.98 31.09 64.50
+Classical (w/ instruct., mean, step 280) 76.18 42.99 85.44 57.63 53.96 82.53 29.94 63.19
+Classical (w/ instruct., last, step 280) 76.54 46.22 86.70 57.79 53.73 82.22 30.13 63.87
+Table 18: Results from all MTEB datasets for finetuning with Mistral-7B.
+Dataset Echo (last) Echo (mean) Classical (last) Classical (mean) Bidirectional (last)
 AmazonCounterfactualClassification 82.97 82.91 80.82 82.21 83.07
@@
 SummEval 30.73 29.46 31.02 30.19 29.32
-Table 16: Results from all MTEB datasets for finetuning with Mistral-7B.